### v1

# Inertial Coordination Games

Abstract
We analyze inertial coordination games: dynamic coordination games with an endogenously changing state that depends on (i) a persistent fundamental that players privately learn about; and (ii) past play. We give a tight characterization of how the speed of learning shapes equilibrium dynamics: the risk-dominant action is selected in the limit if and only if learning is slow such that posterior precisions grow sub-quadratically. This generalizes results from static global games and endows them with an alternate learning foundation. Conversely, when learning is fast, equilibrium dynamics exhibit persistence and limit play is shaped by initial play. Whenever the risk dominant equilibrium is selected, the path of play undergoes a sudden transition when signals are precise, and a gradual transition when signals are noisy.
1   Introduction
Coordination games are at the heart of many economic phenomena. In such games, shocks to the fundamentals can sometimes propagate: as traders attack a currency peg, the central bank has to expand reserves to maintain the peg; as reserves are drained, more traders are incentivized to take advantage of this weakness which, in turn, further drains the central bank’s reserves, enticing even more attacks. But shocks can also fizzle out: traders might learn quite quickly that the central bank’s balance sheet is healthy so that the proportion of the population who are (incorrectly) pessimistic thins out rapidly. Thus, this downward spiral of an initial attack begetting further attacks eventually loses steam. [^1] Similar forces operate with bank runs (via bank liquidity), networked products (via product improvements), firm pricing (via sticky consumers). When do shocks propagate and become self-fulfilling, and when do they fizzle out?
We will show that the speed of private learning is a crucial determinant of whether shocks exhibit persistence. Our model is a dynamic coordination game with an endogenously changing state. A continuum of agents repeatedly decide whether or not to take a risky action whose payoff comprises (i) a persistent fundamental component; and (ii) an endogeneous component which depends on past play. In each period, each agent observes a private signal about the fundamental and makes inferences about the current state, which determines her current behavior. Agents’ current aggregate decisions, in turn, propagate into the future by shaping the evolution of future states. Thus, our environment features both private learning—in which dispersion in beliefs about fundamentals across the population shrinks with time—as well as intertemporal contagion, in which higher current play increases future states which incentivizes higher future play. This endogenous inertia in how the state evolves plays a critical role in the dynamics of play and we call such environments inertial coordination games. [^2] Our model is consistent with extensive empirical documentation of bank failures: Correia, Luck, and Verner (2024) find that “most bank failures are the result of a deterioration in bank solvency.” (in our model, a shock) rather than being purely self-fulfilling. Furthermore, they find that “ Depositors tend to be slow to react to an increased risk of bank failure” (in our model, the intertia).
Our main result (Theorem 1) develops a tight characterization of the relationship between the speed of learning and limit play. We identify a critical threshold for the speed of learning that determines limit play: the risk-dominant action is played in the limit if and only if posterior precisions grow sub-quadratically. Sub-quadratic learning includes canonical learning environments such as (i) independent and identically distributed signals; and (ii) the receipt of a one-time signal. Super-quadratic learning includes social learning, where in each period, agents are randomly matched and sample their partner’s information sets.
Indeed, static global games (Carlsson and Van Damme, 1993) are recovered exactly as a special case of our model in which players only receive a single signal at time $t=1$: since this learning process is sub-quadratic (posterior precisions are unchanged after $t=1$), the risk-dominant profile is—as in global games—played in the limit. This connection arises from observing that best-responding to expected past play along each time step is equivalent in our model to the interim elimination of strictly dominated strategies in the static game. As we make precise in Appendix D, this uncovers a formal connection between “contagion” arguments over type spaces (which encode hierarchies of higher-order beliefs) and “contagion” over time (where the current state depends on past play). Thus, our results generalize the predictions of static global games to sub-quadratic learning and offer a simple dynamic learning foundation for the ‘vanishing noise limit’ commonly invoked in that literature.
A notable implication of our main result is that slow learning implies history independence: the initial shock to aggregate play as well as any finite path of play are irrelevant for limit play. Conversely, when the posterior precision grows super-quadratically, learning is sufficiently fast which induces hysteresis and history-dependence that leads to non-risk dominant limit play. We further characterize all possible equilibria: any monotone selection between initial play-dominance (under common-knowledge of the fundamental state, players best-respond to the conjecture that the proportion of players choosing the risky action is given by initial play) and risk-dominance can be implemented via some appropriate learning rate. Thus, our results offer a simple and unified analysis for when our model’s dynamics exhibit history-dependence such that initial shocks can be self-fulfilling, [^3] See Section I of Krugman (1991) for an excellent discussion of history-dependence and expectations in economics. and when it exhibits history-independence so that limit play is independent of the initial shock.
Finally, our model makes sharp qualitative predictions about the path of play (Proposition 3) whenever learning is slow enough such that the risk-dominant profile is played in the limit. In particular, transition dynamics toward limit play depend starkly on the precision of signals. When signals are precise, aggregate play exhibits a sudden transition from approximately all players playing the non-risk-dominant action to approximately all players playing the risk-dominant action. Conversely, when signals are noisy, aggregate play exhibits a gradual transition as the measure of players playing risk-dominant action increases gradually. These differing transition regimes are driven by differences in time-varying heterogeneity in beliefs. The importance of heterogeneity has been recognized as a crucial determinant of equilibrium uniqueness in coordination games (Morris and Shin, 2006). Our results complement this insight by highlighting that the rate of decrease of heterogeneity driven by learning is also a crucial determinant of the path of play. It also suggests that “spikes” in the time-series of aggregate actions (e.g., withdrawal decisions or adoption of networked goods) can be consistent with transition to limit equilibrium play and need not necessarily correspond to “equilibrium shifts” (Morris and Yildiz (2019)).
Related literature. Our paper directly relates to the large literature on global games. A broad takeaway from the literature on static global games (Carlsson and Van Damme, 1993; Kajii and Morris, 1997; Frankel, Morris, and Pauzner, 2003, among others) is that perturbing common knowledge by introducing a vanishing amount of private uncertainty often selects the risk dominant equilibrium. [^4] But not always (Weinstein and Yildiz, 2007). This approach has been widely deployed in a variety of economic applications to select among multiple equilibria (see, e.g., Morris and Shin (1998); Goldstein and Pauzner (2005) among many others). Our model nests the canonical global games model of Carlsson and Van Damme (1993) with Normal signals as a special case with a private signal at $t=1$ and no learning thereafter. The “contagion argument” in global games where players iteratively eliminate interim dominated strategies is sometimes viewed metaphorically or as a convenient way to select among equilibria. [^5] See, Morris (2002) and Chapter 2 of Spiegler (2024). By developing a simple dynamic environment with endogeneous time-varying states, we thus offer an explicit dynamic foundation for the “contagion” logic of global games. Moreover, we generalize the prediction of risk-dominant selection to the regime of sub-quadratic learning.
Another strand of the literature has investigated equilibrium selection from a non-Bayesian perspective, in particular under exogenous mutations or limited memory (Fudenberg and Harris (1992), Kandori, Mailath, and Rob (1993),Young (1993), Bergin and Lipman (1996), Steiner and Stewart (2008), Block, Fudenberg, and Levine (2019)). In contrast, we propose a model of private Bayesian learning in which the speed of learning and initial conditions jointly mediate limit play. Our finding that super-quadratic learning rates lead limit play to depend on initial conditions is related to the “hysteresis equlibrium” in macroeconomic settings modelled as coordination games discussed in Cooper (1994); Krugman (1991) and, more recently, in Morris and Yildiz (2019). Our results develop a tight characterization of when hysteresis plays a role for limit outcomes, and when initial conditions wash out. [^6] See also Chen and Suen (2016) and Kozlowski, Veldkamp, and Venkateswaran (2020) for a different mechanism of how model uncertainty and rare events interact and give rise to persistence.
Our work also relates to dynamic global games (Angeletos, Hellwig, and Pavan, 2007; Dasgupta, 2007; Dasgupta, Steiner, and Stewart, 2012, among others). In much of this literature, coordination is contemporaneous i.e., time-$t$ play matters for time-$t$ incentives. [^7] Important exceptions are the papers of Dasgupta, Steiner, and Stewart (2012) and Mathevet and Steiner (2013). Our setting is one in which incentives are intertemporally linked because the path of past play shapes the current state which in turn shapes current incentives. In this regard, our paper is perhaps most related to the elegant paper of Mathevet and Steiner (2013) who study finite-time horizon dynamic global games with non-contemporaneous coordination motives where the terminal outcome depends on both the state and path of play. Our models are non-nested and deliver complementary insights. In our model, the state is time-varying and endogenous to past play so players’ payoffs depend on interim play which in turn drives the state-evolution. In Mathevet and Steiner (2013)’s model, ex-post payoffs depend on a binary outcome which is determined by a persistent fundamental and the path of aggregate play. This mapping from paths to a binary outcome allows Mathevet and Steiner (2013) to tractably handle a wide class of payoffs. By contrast, our payoffs are more specific, but evolve endogeneously and take on a continuum of values. We exploit the structure of our environment to derive sharp implications of learning speeds on limit outcomes. [^8] The main result of Mathevet and Steiner (2013) is that aggregate play at the critical state at which outcomes hinge only depends on payoffs; by contrast, we show that limit aggregate play at any state depends sharply on the speed of learning. Note also that Mathevet and Steiner (2013)’s definition of “fast learning” takes the precision of each signal arbitrarily large, which is sufficient for equilibrium existence. This is different from our notion which requires the (weaker) condition of super-quadratic growth in posterior precisions, and where limit equilibrium play always converges.
Finally, our results are closely connected to the literature on learning in coordination games. One interpretation of our model is one of best-response dynamics with Bayesian learning about a fundamental state or past play. To our knowledge, this is novel to the extant literature on learning in games which assumes that past play is perfectly observed and there is no fundamental uncertainty. [^9] See Fudenberg and Levine (1998) for a survey. Our paper is related to an important paper of Crawford (1995) who studies adaptive dynamics in complete information coordination games by modelling player strategies as a linear adjustment rule to be estimated; this is similar to the interpretation of players in our model as best-responding to past play. Thus, a contribution of our paper is to analyze how Bayesian learning about fundamental uncertainty can interact with adaptive learning about strategic uncertainty (Camerer and Hua Ho, 1999). [^10] Camerer and Hua Ho (1999) develop a model of learning in complete information games which features both modes of learning; by contrast, in the best-response interpretation of our model, different modes of learning correspond to different kinds of uncertainty:‘fundamental uncertainty’ is resolved via Bayesian learning but ‘strategic uncertainty’ is resolved via adaptive or reinforcement learning.
2   Model
Payoffs. There is a unit measure of agents indexed by $i\in[0,1]$. [^11] Endow $[0,1]$ with the Lebesgue extension constructed in Sun and Zhang (2009) so that the continuum law of large number holds. Time is discrete and indexed by $t\in\mathcal{T}=\{1,2,\ldots\}$. At each time $t\in\mathcal{T}$, each agent $i$ chooses between a risky action $(a_{it}=1)$ or a safe action $(a_{it}=0)$. Let $\lambda_{t}$ denote the measure of agents who play the risky action at time $t$.
There is an unknown and persistent fundamental $\theta\in\mathbb{R}$. For each $t\in\mathcal{T}$, let $\theta_{t}$ be the state at time $t$ and for each $t>1$, [^12] For any $\alpha\in(0,1)$, analogs of our results apply for $\theta_{t}=\alpha\theta+(1-\alpha)\lambda_{t-1}$; see Appendix E.
$\theta_{t}:=\underbrace{\theta}_{\text{Persistent}}+\underbrace{\lambda_{t-1}}_{\text{Endogenous}}\quad\text{with initial play $\lambda_{0}\in[0,1]$.}$
Each agent $i$’s payoffs at time $t$ are:
$u(a_{it},\theta_{t}):=\begin{cases}\theta_{t}&\text{if }a_{it}=1\\ c&\text{if }a_{it}=0\\ \end{cases}$
$\lambda_{0}$ is common knowledge but interim payoffs as well as $(\lambda_{t})_{t\geq 1}$ are unobserved and inferred via beliefs about the state $\theta$ [^13] An identical assumption is made by Angeletos, Hellwig, and Pavan (2007). Our model is equivalent to one in which agents observe noisy signals of past play $(\lambda_{t})_{t\geq 1}$, or both signals about the state as well as past play—we formalize this an equivalence in Appendix B and discuss this below.
Interpreting initial play. We interpret $\lambda_{0}\in[0,1]$ as a shock. This is equivalent to a one-off and transitory shock to the fundamental or costs at time $1$. [^14] More explicitly, suppose there is zero initial play but there is an initial shock to costs of playing the risky action. Costs are time-varying $c_{1},c_{2},\ldots$ with $c:=c_{2}=c_{3}=c_{4}=\ldots$ and $c_{1}\neq c$. Then, the dynamics of this model with zero initial play is equivalent to the model with $\lambda_{0}=c-c_{1}$. The impact of this shock is equivalent to that of non-zero initial play: positive shocks which incentive the risky action corresponds to high $\lambda_{0}$ and vice versa. In the language of macroeconomics, this corresponds to an unexpected shock (“MIT shock”) and we analyze the dynamics of aggregate play following the shock. In the language of fictitious play, this corresponds to a shock to initial play and we analyze how the speed of learning shapes the limit action profile.
Timing and Information. The timing of the game and the informational environment are as follows. All agents share a common improper uniform prior about $\theta$. At each time $t\in\mathcal{T}$, each agent $i$ receives a private signal $x_{it}\sim N(\theta,\sigma_{t}^{2})$ that is iid across agents and independent across time. Each agent $i$ then updates to her time-$t$ posterior about $\theta$ conditional on the natural filtration $\mathcal{F}_{t}$ generated by past signals $\{x_{is}\}_{s=1}^{t}$. From standard Bayesian updating, this posterior is also Gaussian: $\theta|\mathcal{F}_{t}\sim N(\mu_{it},\eta_{t}^{2})$, where $\mu_{it}:=\eta_{t}^{2}\sum_{s=1}^{t}\sigma_{s}^{-2}x_{is}$ and $\eta_{t}^{2}:=(\sum_{s=1}^{t}\sigma_{s}^{-2})^{-1}$. Finally, each agent $i$ chooses $a_{it}\in\{0,1\}$ to maximize her expected utility. Note that since agents are atomless, they cannot influence future aggregate play, so it is without loss to assume they are myopic. [^15] For example, we could redefine agent $i$’s payoffs as $\max_{(a_{it})_{t\geq 1}}\sum_{t\geq 1}\delta^{t}\mathbb{E}_{t}[u(a_{it},\lambda_{t-1},\theta)]$ for some $\delta\in(0,1)$.
Figure 1: Timing and intertemporal linkages
Limits and Learning. Let the learning process $\Sigma:=(\sigma_{t}^{2})_{t\in\mathcal{T}}$ be a strictly positive sequence of signal variances determining the speed of learning. Note that instead of $\Sigma$, we may equivalently directly specify a nonnegative, increasing sequence of posterior precisions $(\eta_{t}^{-2})_{t}$, since such a sequence is induced by $\sigma_{t}^{-2}:=\eta_{t}^{-2}-\eta_{t-1}^{-2}$. When convenient, we will therefore wlog consider $(\eta_{t}^{-2})_{t}$ as a primitive.
We maintain the following assumption unless otherwise stated.
Assumption 1 (Limit learning).
$\Sigma$ satisfies $\lim_{t\to\infty}\eta_{t}^{-2}=+\infty$.
A key object of analysis in this paper is limit aggregate play at state $\theta$, as a function of signal variances $\Sigma$ and initial aggregate play $\lambda_{0}$. We denote this by $\lambda_{\infty}(\theta|\Sigma,\lambda_{0}):=\lim_{t\to\infty}\lambda_{t}(\theta|\Sigma,\lambda_{0})$. When $(\Sigma,\lambda_{0})$ is understood to be fixed, we omit dependencies. In particular, we study when limit aggregate play converges to the risk dominant (RD) action at state $\theta$, denoted by
$RD(\theta):=\mathbbm{1}\big{\{}\theta\geq c-(1/2)\big{\}}.$
The connection to risk-dominance is as follows: $RD(\theta)$ is the best-response for each agent if she knew $\theta$ and believed that half the population played the risky action. [^17] See Morris and Shin (2003) for a survey of the connection between risk-dominance and continuum player global games. Analogously, define the non-risk dominant action as $NRD(\theta):=1-RD(\theta)$.
It will be helpful to discuss various aspects of our model:
  • Discussion of state transition. Our model is one in which the current state is shaped by past play. For instance, in a currency crisis, if more players attacked the currency in the previous period, this weakens the fundamentals—e.g., the quantity of central bank reserves—which leads to a higher state today. With networked goods, if more players bought the good the previous period, this generates revenue or data for the firm to improve its product which leads to a better product today. We further note that although the current state $\theta_{t}$ only depends directly on $\lambda_{t-1}$, it also depends indirectly on the full (endogeneous) path of $(\lambda_{s})_{s<t-1}$; this is broadly analogous to how an AR$(1)$ process depends indirectly on realizations of the process which are far into the past (although our setting is non-stationary). [^18] An alternate interpretation is that our model is a form of best-response dynamics. Of course, some care is required with this interpretation since past play is unobserved: players are thus best-responding to their best guess of past play. Experimental evidence in (complete information) coordination games find that the vast majority (96%) of play constitute myopic best-responses (Mäs and Nax, 2016).
  • Discussion of Gaussian learning structure. We have explicitly parameterized the model with Gaussian signals; doing so allows us to analyze the law of motion of aggregate play as a discrete dynamical system. It will turn out, however, that our main result (Theorem 1) does not rely on the parametric form of Gaussian learning because we will show that whether risk dominance obtains is a tail event (Corollary 1) and, furthermore, under regularity assumptions, the posterior distribution generated from a sequence of bounded-variance noise structures converges to the normal distribution. [^19] By a straightforward application of the Bernstein–von Mises theorem (Van der Vaart, 2000). Thus, our result characterizing sub/super-quadratic boundary as critical to limit play holds for more general learning processes.
  • Discussion of observability. We have assumed agents do not observe aggregate past play. [^20] Aggregate past play is an injective function of the state, so if agents were to observe this, they would immediately learn the state. However, our model is equivalent to one in which agents receive private signals both about the state, and about aggregate past play. [^21] For instance, see Dasgupta (2007); Trevino (2020). Explicitly, suppose that agent $i\in[0,1]$, in addition to observing $x_{it}\sim N(\theta,\sigma^{2}_{t})$, additionally observes an independent signal distributed $y_{it}\sim N(\Phi^{-1}(\lambda_{t-1}),\tau_{t}^{2})$; this signal structure was introduced by Dasgupta (2007) and used recently by Trevino (2020) to study financial contagion. Then, we can show by induction that equilibrium behavior is equivalent to a model in which players receive a more precise signal about $\theta$ and no information about aggregate past play; Appendix B formalizes this connection. In light of this observation, we load all learning on the state, and agents make inferences about aggregate past play conditional on their beliefs about the state.
  • Discussion of continuum players. We follow the literature on global games by assuming that there is a continuum of atomless players. By the continuum law of large numbers, this implies that conditioned on the fundamental state $\theta$, all aggregate randomness in private signal realizations across the population washes out. This allows us to analyze the dynamics of aggregate play cleanly as the solution to a non-linear deterministic difference equation. In Appendix C we develop a finite-player version of our model with $N<+\infty$ players and show via standard concentration arguments of the empirical distribution of beliefs that limit play behaves similarly in the finite-player version of our model: for all $(\Sigma,\lambda_{0})$ and fundamental $\theta$, $\sup_{N\geq 2}\big{|}\lambda_{\infty}^{N}(\theta)-\lambda_{\infty}(\theta)\big{|}=0$ almost surely (see Theorem 2 in Appendix C).
3   Equilibrium law of motion
We begin by analyzing the joint evolution of aggregate beliefs and aggregate play for the learning process and initial profile of play $(\Sigma,\lambda_{0})$. Appendix A contains the relevant proofs.
Dynamics of threshold beliefs. Recall that agent $i$’s time-$t$ posterior is $N(\mu_{it},\eta_{t}^{2})$. Agent $i$ chooses $a_{it}=1$ at time $t$ if and only if $\mu_{it}\geq\mu_{t}^{*}$, where the belief threshold $\mu_{t}^{*}$ is implicitly defined by the indifference condition
$\mu^{*}_{t}+\mathbb{E}_{\theta\sim N(\mu^{*}_{t},\eta^{2}_{t})}\Big{[}\lambda_{t-1}(\theta)\Big{]}=c$
Let $\Phi$ be the standard Gaussian CDF. Expanding this indifference condition yields the law of motion for belief thresholds $(\mu^{*}_{t})_{t}$, which can be written explicitly as the nonlinear difference equation
$\mu^{*}_{t}=\mu^{*}_{t-1}+\Big{(}\sqrt{\eta^{2}_{t-1}+\eta^{2}_{t}}\Big{)}\cdot\Phi^{-1}(c-\mu^{*}_{t})\quad\text{for all $t\geq 1$}$
with boundary $\mu_{1}^{*}:=c-\lambda_{0}$. Note that the dynamics of $(\mu_{t}^{*})_{t}$ are (i) monotone in $t$; (ii) do not depend on the state $\theta$; and (iii) slow down as agents’ beliefs become more precise. Intuitively, when agents are confident about the state, they are confident that their peers have already coordinated on an action, so the indifference condition should then not vary much.
Dynamics of aggregate play. We now go from threshold beliefs to aggregate play. By the continuum law of large numbers, [^22] See, e.g., Sun and Zhang (2009). the time-$t$ distribution of posterior means in the population of agents is $\mu_{it}\sim N(\theta,\eta_{t}^{2})$. Hence, aggregate time-$t$ play is simply given by
$\lambda_{t}(\theta)=\mathbb{P}_{\mu_{it}\sim N(\theta,\eta_{t}^{2})}\Big{(}\mu_{it}\geq\mu_{t}^{*}\Big{)}=\Phi\left(\frac{\theta-\mu_{t}^{*}}{\eta_{t}}\right)$
Since we assumed limit learning, $\lim_{t\to\infty}\eta_{t}=0$. Taking $t\to\infty$ yields
$\lambda_{\infty}(\theta)=\mathbbm{1}[\theta\geq\mu_{\infty}^{*}]$
Hence, the limit belief threshold $\mu_{\infty}^{*}$ pins down limit aggregate play at all states $\theta$. In particular, $\lambda_{\infty}(\theta)=RD(\theta)$ a.e. if and only if $\mu_{\infty}^{*}=c-(1/2)$. In words, the risky action is adopted by the whole population in the limit if and only if, in the complete information game in which the fundamental $\theta$ is known, playing the risky action is the best response to the uniform conjecture that $1/2$-proportion of players are playing the risky action. Hence, we call $c-(1/2)$ the risk dominance threshold.
We now turn to two simple but informative examples.
Example 1 (Global games).
For $\sigma>0$, let the learning process $\Sigma=(\sigma^{2},\infty,\infty,\ldots)$ be such that agents only receive information at time $t=1$. Since $\eta_{t}^{2}=\sigma^{2}$ for all $t\in\mathcal{T}$, (3) reduces to the stationary difference equation
$\mu_{t}^{*}=\mu_{t-1}^{*}+\sqrt{2}\sigma\Phi^{-1}(c-\mu_{t}^{*})$
For any $\sigma>0$ and $\lambda_{0}\in(0,1)$, this equation has a unique, globally attracting fixed point at $\mu_{\infty}^{*}(\sigma,\lambda_{0})=c-(1/2)$. Hence, for any initial signal precision and initial aggregate play, the risk dominance threshold is attained in the limit. However, without limit learning, limit play is not exactly risk-dominant since some agents remain mistakenly optimistic/pessimistic about the state. But, as the initial signal becomes arbitrarily precise,
$\lim_{\sigma\downarrow 0}\lambda_{\infty}(\theta|\sigma,\lambda_{0})=RD(\theta).$
[…]
Example 2 (Perfect learning).
Consider any learning process $\Sigma$ with $\sigma_{1}=0$, such that $\theta$ is common knowledge from time $t=1$ onwards. Suppose that $\theta+\lambda_{0}>c$. We show by induction that $\lambda_{t}=1$ for all $t\in\mathcal{T}$. For the base case, note that every agent $i$ strictly prefers $a_{it}=1$ if and only if $\theta+\lambda_{0}>c$, which holds by assumption. Hence, $\lambda_{1}=1$ is common knowledge. For the inductive step, assume $\lambda_{t-1}=1$. Then,
$\theta+\lambda_{t-1}-c=\theta+1-c\geq\theta+\lambda_{0}-c>0$
which implies $\lambda_{t}=1$. The argument for $\theta+\lambda_{0}<c$ is analogous so $\lambda_{t}=0$ for all $t\in\mathcal{T}$. Hence, limit play depends on $\lambda_{0}$: $\lambda_{\infty}(\theta|\Sigma,\lambda_{0})=\mathbbm{1}[\theta\geq c-\lambda_{0}].$$\diamondsuit$
Examples 1 and 2 resemble the well-understood gap between small perturbations of common knowledge highlighted in the literature on static coordination games (Example 1). In our environment, complete information does not induce multiplicity because players’ incentives are non-contemporaneous; but it induces history-dependence: limit play is sensitive to initial play. [^23] This is consistent with experimental evidence from coordination games which have largely focused on complete information settings; see, e.g. Van Huyck, Battalio, and Beil (1991).
What are we to make of these differing predictions? In what follows we will provide a unified account of when risk-dominance or history-dependence obtains as a function the learning process. It will turn out that Example 1 is a special case of slow learning since posterior precisions grow sub-quadratically (in Example 1, they do not change beyond $t=1$). On the other hand, Example 2 is a special case of fast learning since posterior precisions are super-quadratic (in Example 2, they “jump immediately to infinity” at time $t=1$). [^24] Translated into a static incomplete information game, this correspond to interspersing interim Bayesian updating with rounds of interim deletion of strictly dominated strategies. Appendix D makes this connection precise.
4   Characterization of limit play
Our main result gives a tight characterization of limit outcomes in terms of the learning rate. All proofs are in Appendix A. We use standard asymptotic notation: $f(t)=O(g(t))$ if there exists constants $\overline{C}<+\infty$ and $T<+\infty$ such that $t\geq T$ implies $f(t)\leq\overline{C}g(t)$. $f(t)=\Omega(g(t))$ if there exists constants $\underline{C}<+\infty$ and $T<+\infty$ such that $t\geq T$ implies $f(t)\geq\underline{C}g(t)$.
Theorem 1.
The following gives a tight characterization of risk-dominance:
  • (i)
    For all $p<2$ and all $\lambda_{0}\in(0,1)$,
[…]
  • (ii)
    For all $q>2$ and all $\lambda_{0}\in(0,1)$,
    $\eta^{-2}_{t}=\Omega(t^{q})\implies$ $\lambda_{\infty}(\theta)=NRD(\theta)$
    for some positive measure interval $(\underline{\theta},\overline{\theta})$.
Theorem 1 identifies the critical threshold for limit play and, in so doing, develops a tight connection between sub-quadratic growth of posterior precisions and risk-dominance. Economically, it suggests that whether or not coordination succeeds, and how it depends on initial play, depends sharply on the learning rate. We discuss the implications of each part in turn.
Theorem 1 (i) generalizes the global games prediction to the case where learning is sufficiently slow. Thus, our model with an endogenously changing state offers an alternate foundation for equilibrium selection. Moreover, it clarifies that sub-quadratic growth of posterior precision is both necessary and sufficient for risk-dominant selection. Thus, although the ‘vanishing noise limit’ in static global games (Carlsson and Van Damme, 1993; Morris and Shin, 2003) is often interpreted as modelling players with vanishing uncertainty, this actually corresponds to slow learning in our dynamic model with endogenously changing states. [^25] We are grateful to Muhamet Yildiz for pointing out this interpretation.
Figure 2: Path of aggregate play under different learning regimes
Parameters: $\theta=0.4$, $\lambda_{0}=0.75$, $c=1$ (not to scale).
(a) Sub-quadratic posterior precision
(b) Super-quadratic posterior precision
Economically, this also implies that shocks to aggregate incentives (which $\lambda_{0}$ can be interpreted as) do not persist when learning is sufficiently slow. Figure 2b (a) illustrates a typical path of aggregate play under slow learning. The true state is $\theta=0.4$ and this governs the location of (heterogeneous) beliefs in the population. The cost of playing the risky action is $c=1$ so under the conjecture that $1/2$-proportion of the population plays the risky action and common-knowledge of the state $\theta=0.4$, the risk-dominant profile is to not play the risky action. Thus, following the initial shock $\lambda_{0}=0.75$, aggregate play converges back to the safe action.
Theorem 1 (ii) breaks risk-dominance by showing that sufficiently fast learning can reintroduce the influence of initial play. This generalizes Example 2’s prediction from instantaneous learning to fast learning. Economically, this implies shocks to aggregate incentives are persistent when learning is fast. Figure 2b (b) illustrates a typical path of aggregate play under super-quadratic learning. Under the conjecture that $\lambda_{0}=0.75$ proportion of players play the risky action and common-knowledge of the state $\theta=0.4$, playing the risky action is indeed a best-response.
The result in Theorem 1 (ii) that sufficiently fast learning in our model can break risk-dominance is related to the idea that public signals can yield non-risk dominant play even as uncertainty about the fundamentals vanish; see Morris and Shin (2003) Chapter 3.2. In both cases, fast private learning or public signals enable players to quickly approach common knowledge of the fundamental and their peers’ play, thus enabling coordination on the non-risk dominant action. Our main result complements and enhances this insight by providing an explicit dynamic foundation, as well as a precise characterization of how fast learning needs to be for this intuition to hold. This allows us to classify several canonical learning environments and their resulting limit outcomes.
Intuition for Theorem 1. Recall our previous observation that the speed of evolution of the time-$t$ indifferent belief $(\mu_{t}^{*})_{t}$ is inversely related to the speed of learning: when learning is slow enough, $(\mu_{t}^{*})_{t}$ gradually converges to the risk dominant threshold $\mu_{\infty}^{*}=c-(1/2)$ from any initial condition. However, when learning is fast enough, the dynamics of $(\mu_{t}^{*})_{t}$ freeze, and consequently $\mu_{\infty}^{*}$ is bounded away from the risk dominant threshold. The proof of Theorem 1 formalizes this intuition by explicitly constructing fictitious threshold sequences that converge to (non)-risk dominant play, matching them to the appropriate learning rates.
Figure 3: Slow learning implies risk dominance
Figure 3 illustrates the forces underlying Theorem 1 (i) for the case where $\lambda_{0}<1/2$. Each panel corresponds to a different time. In each panel, the dotted line represents the state $\theta$, and the red curve is the distribution of posterior means $\mu_{t}$. The blue line is the threshold belief $\mu_{t}^{*}$, and the shaded area is the mass of agents $\lambda_{t}$ who take the risky action because their beliefs exceed the threshold. As time passes, agents learn and $\mu_{t}$ contracts around $\theta$. In Figure 3, learning and hence the speed of this contraction is slow. Consequently, the evolution of $\mu_{t}^{*}$ is relatively fast. As $t\to\infty$, $\mu_{t}^{*}$ converges to $c-(1/2)$ and $\mu_{t}$ converges to a point mass on $\theta$ due to limit learning. Since $\theta\geq c-(1/2)$, aggregate play converges to the risk dominant (risky) action.
Figure 4 illustrates the forces underlying Theorem 1 (ii) for the case where $\lambda_{0}<1/2$. With super-quadratic learning, the speed of contraction of $\mu_{t}$ is fast. Consequently, the evolution of $\mu_{t}^{*}$ is relatively slow. As $t\to\infty$, $\mu_{t}^{*}$ converges to a limit threshold $\mu_{\infty}^{*}$ that is bounded away from $c-(1/2)$, and since $\theta\in(c-(1/2),\mu_{\infty}^{*})$, aggregate play converges to the non-risk dominant (safe) action.
Figure 4: Fast learning implies non-risk dominance
Several corollaries follow relatively straightforwardly from Theorem 1.
Corollary 1 (Risk-dominance is a tail event).
Consider any sub-quadratic learning process $\Sigma=(\sigma_{t}^{2})_{t\in\mathcal{T}}$. Then, for any finite $T<+\infty$ and any $(\sigma^{\prime}_{t})_{t\leq T}$, the risk-dominant action profile is selected almost surely in the limit under the learning process $\Sigma^{\prime}:=[((\sigma^{\prime}_{t})^{2})_{t\leq T},(\sigma_{t}^{2})_{t>T}]$.
This gives an explicitly dynamic foundation for the gap between almost-common knowledge and common knowledge of payoffs in static incomplete information games. Indeed, Corollary 1 suggests that the Gaussian form of private signals is not necessary since, for (i) an arbitrary noise structure (with bounded variance); and (ii) “sufficiently smooth priors”, as $T$ grows large the distribution over posterior beliefs after $T$ grows close to the normal distribution by the Bernstein-von Mises Theorem (Van der Vaart, 2000).
Corollary 2.
The following are limit outcomes of canonical learning processes:
  • (i)
    (IID Gaussian signals) Suppose $\Sigma=(\sigma^{2},\sigma^{2},\ldots)$. Then for any $\lambda_{0}\in(0,1)$ and $\sigma>0$, limit aggregate play is the risk dominant action.
  • (ii)
    (Social learning) At each time $t\geq 2$, player $i$ exchanges information with player $j$. Thus, defining $\Sigma(\sigma)$ to be: $\sigma_{1}^{2}:=\sigma^{2}$ and for all $t\geq 2$, $\sigma_{t}^{2}:=\eta_{t-1}^{2}$, we have that for any $\lambda_{0}$, there exists some $\underline{\sigma}>0$ such that under $\Sigma(\sigma)$ for any $\sigma<\underline{\sigma}$, there exists an open interval of states at which limit aggregate play is the non-risk dominant action.
Corollary 2 (i) shows that limit risk dominant play independent of initial conditions holds for a canonical informational environment. Corollary 2 (ii) studies a social learning setting where at time $1$, each agent observes an iid private signal, and at each subsequent time, each agent is matched at random with another player in the population and observes the other player’s information. This provides a simple informational setting where, if the initial signal is informative enough, non risk-dominant coordination arises.
Comparative statics. Theorem 1 identified a sub- and supercritical regime for how learning speeds impact limit play. We now develop comparative statics on how the set of equilibria vary with primitives. Recall that limit play is jointly pinned down by the limit threshold $\mu^{*}_{\infty}$ and the fundamental $\theta$: $\lambda_{\infty}=\mathbbm{1}(\theta\geq\mu^{*}_{\infty})$. Thus, $\mu^{*}_{\infty}=c-1/2$ corresponds to risk-dominance; as $|\mu^{*}_{\infty}-(c-1/2)|$ becomes large, the non-risk dominant limit profile is selected on a wider set of states.
Proposition 1 (Comparative statics).
The following are comparative statics on $|\mu^{*}_{\infty}-(c-1/2)|$:
  • (i)
    Extreme initial play induces extreme limit play. For any $|\lambda_{0}-(1/2)|\leq|\lambda_{0}^{\prime}-(1/2)|$ and any $\Sigma$,
    $\Big{|}\mu_{\infty}^{*}(\Sigma,\lambda_{0})-(c-1/2)\Big{|}\leq\Big{|}\mu_{\infty}^{*}(\Sigma,\lambda_{0}^{\prime})-(c-1/2)\Big{|}.$
  • (ii)
    Monotonicity in learning speeds. For any $\Sigma\geq\Sigma^{\prime}$,
    $\Big{|}\mu_{\infty}^{*}(\Sigma,\lambda_{0})-(c-1/2)\Big{|}\leq\Big{|}\mu_{\infty}^{*}(\Sigma^{\prime},\lambda_{0})-(c-1/2)\Big{|}.$
  • (iii)
    Early learning rates matter more. Fix any $\Sigma=(\sigma_{t}^{2})_{t}$ and define the pairwise $s\leftrightarrow s^{\prime}$ permutation $\Sigma^{s\leftrightarrow s^{\prime}}:=(\tilde{\sigma}_{t}^{2})_{t}$ where
    $\tilde{\sigma}_{t}^{2}=\begin{cases}\sigma_{t}^{2}&\text{if $t\neq s,s^{\prime}$}\\ \sigma_{s^{\prime}}^{2}&\text{if $t=s$}\\ \sigma_{s}^{2}&\text{if $t=s^{\prime}$}.\end{cases}$
    Then, for any $s\geq s^{\prime}$ and $\sigma_{s}^{2}\geq\sigma_{s^{\prime}}^{2}$,
    $\Big{|}\mu_{\infty}^{*}(\Sigma,\lambda_{0})-(c-1/2)\Big{|}\leq\Big{|}\mu_{\infty}^{*}(\Sigma^{s\leftrightarrow s^{\prime}},\lambda_{0})-(c-1/2)\Big{|}.$
Proposition 1 is intuitive. Part (i) states that the non-risk dominant play expands (i.e., on a wider set of states the non-risk dominant equilibrium is played in the limit) whenever initial play is more extreme on either side of $1/2$; Part (ii) states that non-risk dominant play expands when signals are more precise at each time step (although, per Theorem 1, this only has bite in the superquadratic regime); Part (iii) shows that non-risk dominant play expands when the learning process is permuted so that more informative signals arrive early, and less informative signals arrive late. Intuitively, learning and aggregate behavior at early time steps are more important because they propagate into the future.
Possible limit play. Theorem 1 states that when learning is sufficiently fast, the limit action is non-risk dominant at some open interval of states. Proposition 1 established how this interval varies as primitives change. This raises the following question: what kinds of limit play can be rationalized by some learning rate?
Say that a profile is $\lambda_{0}$-dominant if, under common-knowledge of $\theta$, players play as-if they best-respond to the conjecture that $\lambda_{0}$ proportion of players play the risky action. [^26] This is equivalent to the definition of $p$-dominance in Morris, Rob, and Shin (1995). Our next result establishes that any profile between $\lambda_{0}$-dominance and risk-dominance can be rationalized by some learning speed.
Proposition 2 (The image of learning rates).
For any $\lambda_{0}\in(0,1)$,
$\bigcup_{\Sigma}\lambda_{\infty}(\theta|\Sigma,\lambda_{0})=\bigcup_{\begin{subarray}{c}\text{$\mu^{*}$ between}\\ \text{$c-\lambda_{0}$ and $c-1/2$}\end{subarray}}\mathbbm{1}(\theta\geq\mu^{*}).$
In particular, for any threshold $\mu^{*}$ between $c-\lambda_{0}$ and $c-(1/2)$, there exists some learning rate $\Sigma\geq 0$ such that:
[…]
The intuition behind Proposition 2 is that for any time $t$, we can arbitrarily restrict the evolution of the system by choosing a sufficiently precise signal for time $t$. This verifies that any level of coordination on the non-risk dominant action can be rationalized by some sequence of signal noises. [^27] Recall that $\mu_{1}^{*}=c-\lambda_{0}$ and $\mu_{t}^{*}$ evolves monotonically away. Together with our comparative static on learning speeds (Proposition 1 (ii)) this also implies $\lim_{\Sigma\downarrow 0}\lambda_{\infty}(\theta|\Sigma,\lambda_{0})=\mathbbm{1}(\theta\geq c-\lambda_{0})$—that is, in the vanishing noise limit along the whole path, limit play converges to that of the complete environment with all players best-responding to $\lambda_{0}$.
Proposition 2 also highlights the following asymmetry between initial play and learning rates: when initial play is large ($\lambda_{0}>1/2$ large), slow learning hurts efficiency since the set of fundamental states at which the payoff-dominant action is played shrinks to $\{\theta\geq c-1/2\}$; conversely, when initial play is small ($\lambda_{0}<1/2$ small), slow learning improves efficiency since on the set of fundamentals $\{\theta\geq c-1/2\}$ the payoff-dominant action is played in the limit.
5   Transition dynamics
We have thus far focused on limit play. We now turn to studying the connection between the speed of learning and the time path of aggregate play. We will consider the iid signal case such that the learning process $\Sigma(\sigma)$ is parametrized by the standard deviation of each period’s signal $\sigma>0$. We focus on the iid case because it is a canonical model of learning, and makes results easier to state since $\Sigma$ is now parametrized by a scalar; nonetheless, the qualitative features of our results extend beyond this case. Within this iid learning environment, recall that Corollary 2 (i) implies that the limit action is risk dominant for all fundamentals and initial conditions. The following result shows that the path of aggregate play exhibits qualitatively different behavior depending on learning rates.
Proposition 3.
Fix any $\lambda_{0}\in(0,1)$ and $\theta$ between $c-1/2$ and $c-\lambda_{0}$. Define $\overline{\beta}:=1+|c-(1/2)-\theta|/2$. For any $\epsilon>0$, $\alpha\in(0,1)$, and $\beta\in(1,\overline{\beta})$, there exists $\underline{\sigma},\overline{\sigma}>0$ such that:
  • (i)
    Sudden transition. For all $\sigma<\underline{\sigma}$, there exists $T(\sigma)\geq 1$ such that: for all $1\leq t\leq\alpha T(\sigma)$, $|\lambda_{t}(\theta)-NRD(\theta)|<\epsilon$ and for all $t\geq\beta T(\sigma)$, $|\lambda_{t}(\theta)-RD(\theta)|<\epsilon$;
  • (ii)
    Gradual transition. For all $\sigma>\overline{\sigma}$ and all $t\geq 1$, $|\lambda_{t+1}(\theta)-\lambda_{t}(\theta)|<\epsilon$.
Proposition 3 says that, within the regime of iid learning, sufficiently fast learning yields a sudden transition from aggregate non-risk dominant to risk dominant play that occurs around some finite time. Conversely, sufficiently slow learning yields sufficiently gradual transitions in aggregate play.
Figure 5: Mechanism underlying transition regimes
The intuition behind Proposition 3 is illustrated in Figure 5. In the case where learning is slow, the distribution of posterior means is dispersed so as the threshold changes, a relatively small mass of new agents take the risk dominant action. Conversely, when learning is fast, the distribution of posterior means contracts around $\theta$ so that there is a sudden spike in agents taking the risk dominant action when it is crossed.
Figure 6b numerically illustrates the time path of play $(\lambda_{t})_{t}$ for iid learning. Panel (a) shows the dynamics when $\sigma=1$; panel (b) shows the dynamics when $\sigma=0.01$. In particular, when learning is fast, an analyst would observe next-to no change in the measure of agents playing the risky action before a sudden spike.
Figure 6: Gradual vs sudden transition to risk dominant play
Parameters: $\theta=0.6$, $\lambda_{0}=0.2$, $c=1$.
(a) $\sigma=1$
(b) $\sigma=0.01$
Proposition 3 is subtly connected to the literature on coordination games. A common theme in this literature is that heterogeneity—in preferences, beliefs, etc.—across players can induce uniqueness in static supermodular games (Morris and Shin, 2006). A dynamic version of this insight is that with lots of hetrogeneity, as aggregate incentives shift, the set of equilibria vary smoothly—this is analogous to Part (i) of Proposition 3; by contrast, with little heterogeneity, the set of equilbiria can vary discontinuously—this is analogous to Part (ii) of Proposition 3. The key insight of Proposition 3 is that learning rates induce time-varying hetrogeneity in beliefs which, in turn, matters for transition dynamics. [^28] Koh and Morris (2022) relates heterogeneity to the speed of contagion in a networked population. This can be equivalently interpreted (see, e.g., Morris (1997)) as best-response dynamics or interim deletion of strictly dominated strategies where each player’s beliefs and higher-order beliefs are encoded by the network. However, Koh and Morris (2022) study speed of contagion for graphs/beliefs which are held fixed i.e., in the incomplete information game interpretation, there is no learning.
Economically, Proposition 3 offers a simple alternate interpretation of sudden changes in the time-series of aggregate play: when learning is precise, players have very similar beliefs by the time the “tipping point” arrives and the state has increased to the point that is optimal for such players to play the risky action. Sudden spikes have been commonly interpreted as “equilibrium shifts” driven by shocks or new public information (Chwe, 2013; Morris and Yildiz, 2019) interacting with uniform rank beliefs. Our result offers an alternate explanation driven by private learning and time-varying heterogeneity.
6   Concluding Remarks
We have developed a simple but novel model of dynamic coordination with (i) Bayesian learning about fundamental uncertainty; and (ii) adaptive best-response dynamics. This has allowed us to combine and unify insights from the literature on best-response dynamics in complete-information coordination games (Crawford, 1995) and global games (Carlsson and Van Damme, 1993; Morris and Shin, 2003). Our model serves as a natural description of many economic environments with both shrinking uncertainty about fundamentals (e.g., solvency of a bank, strength of the economy, quality of a networked product) as well as an endogenous time-varying state which is influenced by aggregate past play.
Within our environment, we uncovered a tight connection between learning rates and limit play (Theorem 1) and established sub/super-quadratic reduction in posterior variance as the critical threshold at which limit play undergoes a shift away from risk-dominance. We also show that, subject to initial conditions, any shift from risk-dominance can be attained by some learning rate. Moreover, in the sub-quadratic regime, we further identified qualitatively different behavior in the path of aggregate play (Proposition 3): when learning is slow, transition dynamics are gradual; when learning is fast, sudden.
APPENDIX TO INERTIAL COORDINATION GAMES
Appendix A Omitted proofs
Section A.1 contains several useful lemmas that characterize the dynamics of our model. Sections A.2-A.5 collect all remaining proofs. We sometimes assume $\lambda_{0}>\frac{1}{2}$ without loss of generality because symmetric arguments apply to $\lambda_{0}<\frac{1}{2}$.
A.1  System Dynamics
Lemmas 1-3 show that in equilibrium, agents use a cutoff strategy with respect to posterior means. Lemmas 4-7 study useful auxiliary system dynamics.
Lemma 1.
At each time $t\geq 1$ and for each player $i$, $a_{it}=1$ if and only if $\mu_{it}\geq\mu_{t}^{*}$, where $\mu_{1}^{*}:=c-\lambda_{0}$ and for all $t>1$, $\mu_{t}^{*}$ is the unique solution to the nonlinear difference equation
$\mu_{t}^{*}+\Phi\left(\frac{\mu_{t}^{*}-\mu_{t-1}^{*}}{\sqrt{\eta_{t-1}^{2}+\eta_{t}^{2}}}\right)=c$ ($\mu^{*}$-LoM)
Hence, for each $t\geq 1$,
$\lambda_{t}(\theta)=\Phi\left(\frac{\theta-\mu_{t}^{*}}{\eta_{t}}\right)$ ($\lambda$-LoM)
Proof of Lemma 1.
We prove Lemma 1 by induction. At $t=1$, each agent $i$’s posterior belief is given by $\theta|\mathcal{F}_{i1}\sim N(x_{i1},\sigma_{1}^{2})$. Hence, she takes $a_{i1}=1$ if and only if $\mathbb{E}_{i1}[\theta_{1}]=\mu_{i1}=x_{i1}\geq c-\lambda_{0}=\mu_{1}^{*}$. Since $x_{i1}|\theta\sim N(\theta,\sigma_{1}^{2})$,
$\lambda_{1}(\theta)=\mathbb{P}_{\theta}(x_{i1}\geq\mu_{1}^{*})=1-\Phi\left(\frac{\mu_{1}^{*}-\theta}{\sigma_{1}}\right)=\Phi\left(\frac{\theta-\mu_{1}^{*}}{\sigma_{1}}\right).$
Hence, ($\lambda$-LoM) holds for $t=1$.
Now, for $t>1$, suppose that for all agents $j$, $a_{j,t-1}=1$ if and only if $\mu_{j,t-1}\geq\mu_{t-1}^{*}$. By standard Gaussian-Gaussian Bayesian updating, we have $\theta|\mathcal{F}_{it}\sim N(\mu_{it},\eta_{t}^{2})$, where
$\mu_{it}=\sum_{s=1}^{t}\frac{\sigma_{s}^{-2}}{\eta_{t}^{-2}}x_{is}\quad\text{and}\quad\eta_{t}^{2}=\left(\sum_{s=1}^{t}\sigma_{s}^{-2}\right)^{-1}.$
Conditional on $\theta$, the time-$(t-1)$ distribution of posterior means is $\mu_{j,t-1}|\theta\sim N(\theta,\eta_{t-1}^{2})$. Since agents’ information sets are independent conditional on $\theta$, this implies that, under agent $i$’s time-$t$ posterior belief, the marginal distribution of time-$(t-1)$ posterior means is $\mu_{j,t-1}|\mathcal{F}_{it}\sim N(\mu_{it},\eta_{t}^{2}+\eta_{t-1}^{2})$. Hence, by the inductive step, $i$ takes $a_{it}=1$ if and only if
$c\leq\mathbb{E}_{it}\left[\theta+\lambda_{t-1}(\theta)\right]=\mu_{it}+\mathbb{P}_{it}(\mu_{j,t-1}\geq\mu_{t-1}^{*})=\mu_{it}+\Phi\left(\frac{\mu_{it}-\mu_{t-1}^{*}}{\sqrt{\eta_{t-1}^{2}+\eta_{t}^{2}}}\right)$
Let $F(\mu_{it})$ be the RHS function. Since $F$ is continuous and strictly increasing in $\mu_{it}$, and $F(c)>c$ and $F(c-1)<c$ hold, $F(\mu_{it})=c$ has a unique solution $\mu_{t}^{*}\in(c-1,c)$. Hence, there exists a unique cutoff $\mu_{t}^{*}$, defined implicitly by ($\mu^{*}$-LoM) at time $t$, such that $i$ takes $a_{it}=1$ if and only if $\mu_{it}\geq\mu_{t}^{*}$.
Finally, since $\mu_{it}|\theta\sim N(\theta,\eta_{t}^{2})$, we see that
$\lambda_{t}(\theta)=\mathbb{P}_{\theta}(\mu_{it}\geq\mu_{t}^{*})=\Phi\left(\frac{\theta-\mu_{t}^{*}}{\eta_{t}}\right)$
and hence ($\lambda$-LoM) holds at time $t$, as desired.
Lemma 2.
$(\mu_{t}^{*})_{t}$ has a unique steady state $\mu_{ss}^{*}=c-1/2$. For any $t\in\mathcal{T}$, $\mu_{t+1}^{*}>\mu_{t}^{*}$ if and only if $\mu_{t}^{*}<\mu_{ss}^{*}$.
Proof.
Assume $\mu_{t-1}^{*}=c-(1/2)$. Since $\mu_{t}^{*}:=c-(1/2)$ solves ($\mu^{*}$-LoM) and it has a unique solution, $c-(1/2)$ is a steady state. To see that it’s the unique steady state, assume $\mu_{t-1}^{*}=\mu_{t}^{*}$. Then, $\mu_{t}^{*}=c-(1/2)$, as desired.
Next, fix any $t\in\mathcal{T}$. By ($\mu^{*}$-LoM), we have
$\mu_{t+1}^{*}>\mu_{t}^{*}$
[…]
Hence, if $\mu_{t+1}^{*}>\mu_{t}^{*}$, $\mu_{t}^{*}<\mu_{t+1}^{*}<\mu_{ss}^{*}$. Conversely, if $\mu_{t+1}^{*}\leq\mu_{t}^{*}$, which is equivalent to $\mu_{t+1}^{*}\geq\mu_{ss}^{*}$, we have $\mu_{t}^{*}\geq\mu_{t+1}^{*}\geq\mu_{ss}^{*}$. ∎
Lemma 3.
  • (i)
    If $\lambda_{0}>\frac{1}{2}$, then $\mu_{t}^{*}\uparrow\mu_{\infty}^{*}\in(\mu_{1}^{*},\mu_{ss}^{*}]$.
  • (ii)
    If $\lambda_{0}<\frac{1}{2}$, then $\mu_{t}^{*}\downarrow\mu_{\infty}^{*}\in[\mu_{ss}^{*},\mu_{1}^{*})$.
  • (iii)
    $|\mu^{*}_{t+1}-\mu_{t}^{*}|$ is decreasing in $t$ [^29] Note that parts (i),(ii), and (iii) of this Lemma hold without Assumption 1.
  • (iv)
    $\lambda_{\infty}(\theta)=\mathbbm{1}[\theta\geq\mu_{\infty}^{*}]$ a.e.
Proof.
Parts (i) and (ii). Wlog, assume $\lambda_{0}>\frac{1}{2}$. Since Lemma 2 implies that $(\mu_{t}^{*})_{t}$ is a strictly increasing sequence bounded above by $\mu_{ss}^{*}$, it converges to $\mu_{\infty}^{*}\in(\mu_{1}^{*},\mu_{ss}^{*}]$.
Part (iii). Note that ($\mu^{*}$-LoM) implies
$|\mu^{*}_{t+1}-\mu_{t}^{*}|=\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}\left|\Phi^{-1}(c-\mu_{t+1}^{*})\right|.$
Note that $\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}$ is non-increasing in $t$. Wlog, assume $\lambda_{0}>\frac{1}{2}$. Then, $(c-\mu_{t+1}^{*})$ decreasing in $t$ implies $\Phi^{-1}(c-\mu_{t+1}^{*})>0$ is decreasing in $t$, and therefore $|\Phi^{-1}(c-\mu_{t+1}^{*})|$ is decreasing in $t$. Hence, $|\mu^{*}_{t+1}-\mu_{t}^{*}|$ is also decreasing in $t$.
Part (iv). Wlog, assume $\lambda_{0}>\frac{1}{2}$. First, we show $(\theta-\mu_{t}^{*})/\eta_{t}\to-\infty$ when $\theta<\mu_{\infty}^{*}$. Since $((\theta-\mu_{t}^{*})/\eta_{t})_{t}$ is decreasing and becomes negative for sufficiently large $t$, it should either converge to some negative constant or diverge to $\infty$- ∞. Since $\lim_{t\to\infty}(\theta-\mu_{t}^{*})<0$, $\lim_{t\to\infty}(\theta-\mu_{t}^{*})/\eta_{t}=-\infty$ must hold by limit learning.
Next, we show $(\theta-\mu_{t}^{*})/\eta_{t}\to\infty$ when $\theta>\mu_{\infty}^{*}$. Lemma 3 (i) implies
$0<\frac{\theta-\mu_{\infty}^{*}}{\eta_{t}}<\frac{\theta-\mu_{t}^{*}}{\eta_{t}}.$
Since $(\theta-\mu_{\infty}^{*})/\eta_{t}\to\infty$ as $t\to\infty$, $(\theta-\mu_{t}^{*})/\eta_{t}\to\infty$ holds. ∎
Lemma 4.
Define $\gamma_{1}:=\Phi^{-1}(\lambda_{0})$, and for $t\geq 2$, define
$\gamma_{t}:=\frac{\mu_{t}^{*}-\mu_{t-1}^{*}}{\sqrt{\eta_{t}^{2}+\eta_{t-1}^{2}}}\quad\text{and}\quad A_{t}:=\sqrt{\eta_{t}^{2}+\eta_{t-1}^{2}}>0$
Then,
  • (i)
    $(\gamma_{t})_{t\geq 2}$ obeys the law of motion
    $A_{t}\gamma_{t}=\Phi(\gamma_{t-1})-\Phi(\gamma_{t})$ ($\gamma$-LoM)
  • (ii)
    If $\lambda_{0}>(1/2)$, then $\gamma_{t}\downarrow\gamma_{\infty}\geq 0$.
  • (iii)
    If $\lambda_{0}<(1/2)$, then $\gamma_{t}\uparrow\gamma_{\infty}\leq 0$.
Proof.
Part (i). Subtracting ($\mu^{*}$-LoM) at time $t-1$ from ($\mu^{*}$-LoM) at time $t$ yields:
$A_{t}\gamma_{t}=\mu_{t}^{*}-\mu_{t-1}^{*}=\Big{(}c-\Phi(\gamma_{t})\Big{)}-\Big{(}c-\Phi(\gamma_{t-1})\Big{)}=\Phi(\gamma_{t-1})-\Phi(\gamma_{t})$
Parts (ii) and (iii). Given the initial condition $\mu_{1}^{*}=c-\lambda_{0}$, defining $\gamma_{1}:=\Phi^{-1}(\lambda_{0})$ yields $\mu_{1}^{*}=c-\Phi(\gamma_{1})$. Note that for all $t\geq 2$,
$\gamma_{t-1}\geq\gamma_{t}\iff\Phi(\gamma_{t-1})\geq\Phi(\gamma_{t})\iff\gamma_{t}\geq 0$
Wlog, assume If $\lambda_{0}>(1/2)$, then $\gamma_{1}>0$. Suppose for a contradiction that $\gamma_{2}<0$: then $\gamma_{1}<\gamma_{2}<0$, which is a contradiction since $\gamma_{1}>0$ by definition. Hence, $\gamma_{1}\geq\gamma_{2}\geq 0$. By induction, $(\gamma_{t})_{t}$ is a monotone decreasing sequence bounded below by 00, so it converges: $\gamma_{t}\downarrow\gamma_{\infty}\geq 0$, as desired. ∎
Lemma 5.
For any $A>0$, let $F_{A}:[0,\infty)\rightarrow[0,\infty)$ be implicitly defined as follows:
$AF_{A}(x)=\Phi(x)-\Phi(F_{A}(x))$
Then,
  • (i)
    Fix any $x>0$. $F_{A}(x)$ is decreasing in $A$, and $F_{A}(x)\downarrow 0$ as $A\to\infty$.
  • (ii)
    Fix any $A>0$. $F_{A}(x)$ is increasing in $x$.
Proof.
Part (i). Fix any $x>0$. For any $A^{\prime}>A>0$, note that
$A^{\prime}F_{A^{\prime}}(x)-AF_{A}(x)=\Phi(F_{A}(x))-\Phi(F_{A^{\prime}}(x))$
Suppose for a contradiction that $F_{A^{\prime}}(x)\geq F_{A}(x)$. Since $\Phi$ is increasing, the above implies
$A^{\prime}F_{A^{\prime}}(x)\leq AF_{A}(x)$
which is a contradiction since $A^{\prime}>A$.
By an exactly analogous argument as in Lemma 4(ii), note that $x>0$ implies $x\geq F_{A}(x)\geq 0$, which implies $\Phi(x)-\Phi(F_{A}(x))\in[0,\Phi(x)-(1/2)]$ and hence
$F_{A}(x)\in\Big{[}0,A^{-1}(\Phi(x)-(1/2))\Big{]}\Rightarrow F_{A}(x)\downarrow 0\quad\text{as }A\to\infty$
as desired.
Part (ii). Let $y=F_{A}(x)$ and use the implicit function theorem to differentiate it with respect to $x$:
$Ay^{\prime}=\varphi(x)-\varphi(y)y^{\prime}$
$\implies y^{\prime}=\frac{\varphi(x)}{A+\varphi(y)}>0,$
where $\varphi$ is a standard normal PDF. ∎
Note that ($\gamma$-LoM) states: $\gamma_{t}=F_{A_{t}}(\gamma_{t-1})$.
Lemma 6.
For any $\gamma^{*}>0$, let $A_{\gamma^{*}}:[\gamma^{*},\infty)\rightarrow[0,\infty)$ be defined as:
$A_{\gamma^{*}}(x):=\frac{\Phi(x)-\Phi\Big{(}(\gamma^{*}+x)/2\Big{)}}{(\gamma^{*}+x)/2}$
Then, there exists $m>\gamma^{*}$ such that $A_{\gamma^{*}}(x)>0$ and $A_{\gamma^{*}}^{\prime}(x)>0$ for all $x\in(\gamma^{*},m)$.
Proof.
Note that
$A_{\gamma^{*}}^{\prime}(x)=\left(\frac{\gamma^{*}+x}{2}\right)^{-2}\left(\frac{\gamma^{*}+x}{2}\left[\varphi(x)-\frac{1}{2}\varphi\left(\frac{\gamma^{*}+x}{2}\right)\right]-\frac{1}{2}\left[\Phi(x)-\Phi\left(\frac{\gamma^{*}+x}{2}\right)\right]\right),$
and
$A_{\gamma^{*}}^{\prime}(\gamma^{*})=\frac{\gamma^{*}\varphi(\gamma^{*})}{2(\gamma^{*})^{2}}>0.$
Since $A_{\gamma^{*}}^{\prime}$ is continuous, there exists $m>\gamma^{*}$ such that $A_{\gamma^{*}}^{\prime}(x)>0$ for all $x\in[\gamma^{*},m)$. Since $A_{\gamma^{*}}(\gamma^{*})=0$ holds by definition, this implies $A_{\gamma^{*}}(x)>0$ for all $x\in[\gamma^{*},m)$. ∎
Note that ($\gamma$-LoM) states: $F_{A_{\gamma^{*}}(\gamma_{t-1})}(\gamma_{t-1})=(\gamma^{*}+\gamma_{t-1})/2$.
Lemma 7.
For any $\gamma_{1}>0$, let $A_{\gamma_{1}}:[0,\gamma_{1}]\rightarrow[0,\infty)$ be defined as
$A_{\gamma_{1}}(x):=\frac{\Phi(\gamma_{1})-\Phi(x)}{x}.$
Then, $A_{\gamma_{1}}(\gamma_{1})=0$ and $A_{\gamma_{1}}$ is decreasing on $[0,\gamma_{1}]$. Thus, for any fixed $\gamma^{*}\in(0,\gamma_{1})$,
$A_{\gamma_{1}}(x)\uparrow A_{\gamma_{1}}(\gamma^{*})=:\bar{A}^{*}$
holds as $x\downarrow\gamma^{*}$.
Proof.
Note that
$A_{\gamma_{1}}^{\prime}(x)=x^{-2}\Big{[}-x\varphi(x)-\Big{(}\Phi(\gamma_{1})-\Phi(x)\Big{)}\Big{]}<0\quad\forall x\in[0,\gamma_{1}]$
Then $A_{\gamma_{1}}(x)\uparrow\bar{A}^{*}$ follows from continuity of $A_{\gamma_{1}}(\cdot)$.
Note that ($\gamma$-LoM) states: $F_{A_{\gamma_{1}}(x)}(\gamma_{1})=x$. With these lemmas in hand, we can understand how $(\gamma_{t})_{t}$ changes as a function of $(A_{t})_{t}$.
A.2  Theorem 1 and Corollary 2
Without loss of generality, assume $\lambda_{0}>\frac{1}{2}$.
Proof of Theorem 1 (i).
Suppose that $\eta_{t}^{-2}\leq Ct^{p}$ holds for all $t\geq T$. Then we have
$\eta_{t}^{-2}\leq Ct^{p}\implies\eta_{t}^{2}\geq C^{-1}t^{-p}\implies\eta_{t}^{2}+\eta_{t-1}^{2}\geq 2\eta_{t}^{2}\geq 2C^{-1}t^{-p}$
$\implies(\eta_{t}^{2}+\eta_{t-1}^{2})^{-1/2}\leq(C/2)^{1/2}t^{p/2}$
for $t\geq T$. By Lemma 3 (iii), $\mu_{t+1}^{*}-\mu_{t}^{*}$ is decreasing in $t$, which implies
$\mu_{t+1}^{*}-\mu_{t}^{*}\leq t^{-1}\sum_{s=1}^{t}(\mu_{s+1}^{*}-\mu_{s}^{*})=t^{-1}(\mu_{t+1}^{*}-\mu_{1}^{*})\leq t^{-1}.$
Using this inequality, we can apply the squeeze theorem and compute $\mu_{\infty}^{*}$ as follows:
$0\leq\frac{\mu_{t+1}^{*}-\mu_{t}^{*}}{\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}}\leq\underbrace{(C/2)^{1/2}t^{p/2-1}}_{\to 0\,\,\,\,\because\,p<2}$
$\implies\mu_{\infty}^{*}=c-\Phi\left(\lim_{t\to\infty}\frac{\mu_{t+1}^{*}-\mu_{t}^{*}}{\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}}\right)=c-(1/2)$
and the conclusion follows from Lemma 3 (iv). ∎
Proof of Theorem 1 (ii).
Throughout this proof, fix any $q>2$ and (wlog) $\lambda_{0}>1/2$. When convenient, we omit dependencies on these variables. Let “p.m.” denote “on a set of positive (Lebesgue) measure.”
Step 1. First, we show the following lemma.
Lemma 8.
There exists $\underline{C}>0$ such that, for any $T\geq 1$,
[…]
Proof of Lemma 8.
Let $\gamma_{1}:=\Phi^{-1}(\lambda_{0})$ and define the following expressions:
[…]
$K:=\left[\frac{1}{2}\frac{\varphi(\gamma_{1})}{S}\right]^{2},\ \ \ C:=LK,\ \ \ \underline{C}:=C^{-1}$
Note that $\underline{C}$ does not depend on $t$. We begin with the case $T=1$. Fix any $\Sigma$ with induced posterior precisions $(\eta_{t}^{-2})_{t}$ such that $\eta_{t}^{-2}\geq\underline{C}t^{q}$ for all $t\geq 1$. Then the following implications hold:
$\eta_{t}^{-2}\geq\underline{C}t^{q}\quad\forall t\geq 1$
$\implies\eta_{t}^{2}\leq Ct^{-q}\quad\forall t\geq 1$
$\implies\eta_{t}^{2}+\eta_{t-1}^{2}\leq C(t^{-q}+(t-1)^{-q})\quad\forall t\geq 2$
$\implies(A_{t})^{2}:=\eta_{t}^{2}+\eta_{t-1}^{2}\leq Kt^{-q}\quad\forall t\geq 2$
$\implies A_{t}\leq\frac{1}{2}\frac{\varphi(\gamma_{1})}{S}t^{-r}\quad\forall t\geq 2$
where the third implication follows from
$\frac{C}{K}=L=\frac{2^{-q}}{2^{-q}+1}\leq\frac{t^{-q}}{t^{-q}+(t-1)^{-q}}\quad\forall\ t\geq 2$
since
$\text{sign}(L^{\prime}(t))=\text{sign}\Big{(}(t-1)^{-1}-t^{-1}\Big{)}=1.$
Next, for each $\epsilon\in(0,\gamma_{1})$ and each $t\geq 2$, define:
$\Delta_{t}(\epsilon):=\frac{\gamma_{1}-\epsilon}{S}t^{-r},\ \ \ \gamma_{t}(\epsilon):=\gamma_{1}-\sum_{s=2}^{t}\Delta_{t}(\epsilon),\ \ \ A_{t}(\epsilon):=\frac{\Phi(\gamma_{t-1}(\epsilon))-\Phi(\gamma_{t}(\epsilon))}{\gamma_{t}(\epsilon)}$
with the initial condition $\gamma_{1}(\epsilon)=\gamma_{1}$. Observe that $\gamma_{t}(\epsilon)\downarrow\epsilon$ as $t\to\infty$ and
$\Phi(\gamma_{t-1}(\epsilon))-\Phi(\gamma_{t}(\epsilon))\geq\varphi(\gamma_{t-1}(\epsilon))(\gamma_{t-1}(\epsilon)-\gamma_{t}(\epsilon))$
$\geq\varphi(\gamma_{1})\Delta_{t}(\epsilon)\geq\frac{\gamma_{t}(\epsilon)}{\gamma_{1}}\varphi(\gamma_{1})\Delta_{t}(\epsilon)\quad\forall t\geq 2$
$\implies\frac{\varphi(\gamma_{1})}{\gamma_{1}}\Delta_{t}(\epsilon)\leq A_{t}(\epsilon)\quad\forall t\geq 2$
because $(\gamma_{t}(\epsilon))_{t}$ is decreasing, and $\gamma_{t}(\epsilon)<\gamma_{1}$ holds for all $t\geq 2$. Also note that
$\lim_{\epsilon\downarrow 0}\frac{\varphi(\gamma_{1})}{\gamma_{1}}\frac{\gamma_{1}-\epsilon}{S}=\frac{\varphi(\gamma_{1})}{S}$
Choose $\epsilon>0$ small enough, independent of $t$, such that
$A_{t}\leq\frac{1}{2}\frac{\varphi(\gamma_{1})}{S}t^{-r}\leq\frac{\varphi(\gamma_{1})}{\gamma_{1}}\frac{\gamma_{1}-\epsilon}{S}t^{-r}\leq A_{t}(\epsilon)\quad\ \forall\ t\geq 2.$
By Lemma 5, we can show $\gamma_{t}\geq\gamma_{t}(\epsilon)$ for $t\geq 2$ inductively. Hence,
$\gamma_{\infty}=\lim_{t\to\infty}F_{A_{t}}(\gamma_{t-1})\geq\lim_{t\to\infty}F_{A_{t}(\epsilon)}(\gamma_{t-1}(\epsilon))=\epsilon>0$
where $\gamma_{\infty}$ is the limit of $(\gamma_{t})_{t}$ induced by $\Sigma$.
Next, fix any $T\geq 1$, and let $\underline{C}$ be the constant derived above. Fix any $\Sigma$ satisfying $\eta_{t}^{-2}\geq\underline{C}t^{q}$ for all $t\geq T$, and define:
$\tilde{\underline{C}}(\Sigma):=L^{-1}\left[\frac{1}{2}\frac{\varphi(\gamma_{T})}{S}\right]^{-2}\leq\underline{C}$
where the inequality follows because $\gamma_{T}\leq\gamma_{1}$. Hence, $\Sigma$ satisfies
$\eta_{\tilde{t}}^{-2}\geq\underline{C}\{\tilde{t}+(T-1)\}^{q}\geq\underline{C}(\tilde{t})^{q}\geq\tilde{\underline{C}}(\Sigma)(\tilde{t})^{q}\quad\forall\tilde{t}\geq 1$
where $\tilde{t}:=t-(T-1)$. By an exactly analogous argument as the $T=1$ case, $\gamma_{\infty}>0$ holds as desired. ∎
Step 2. To finish the argument, we want to show:
[…]
Fix any $C>0$, $T\geq 1$, and $\Sigma$. Fix any $q^{\prime}\in(2,q)$, and choose $T^{\prime}\geq T$ large enough such that
$C(T^{\prime})^{q-q^{\prime}}\geq\underline{C}(q^{\prime},\lambda_{0}),$
where $\underline{C}(q^{\prime},\lambda_{0})$ is the constant from Lemma 8. By Lemma 8, the following implications hold:
$\eta_{t}^{-2}\geq Ct^{q}=Ct^{q-q^{\prime}}\cdot t^{q^{\prime}}\quad\forall t\geq T$
$\implies\eta_{t}^{-2}\geq\underline{C}(q^{\prime},\lambda_{0})t^{q^{\prime}}\quad\forall t\geq T^{\prime}$
[…]
Proof of Corollary 2.
Part (i). If $\sigma_{t}=\sigma$ for all $t\geq 1$, then $\eta_{t}^{-2}=\sigma^{-2}t$. Take $C:=\sigma^{-2}$ and $p:=1$, and the conclusion directly follows from Theorem 1 (i).
Part (ii). If $\sigma_{1}:=\sigma$ and $\sigma_{t}:=\eta_{t-1}$ for all $t\geq 2$, then
$\eta_{t}^{-2}=\sigma^{-2}2^{t-1}$
Fix $q=3$ and any $\lambda_{0}$, and let $\underline{C}$ be the constant from Lemma 8 when $q=2$. Choose $\underline{\sigma}>0$ small enough such that
$\underline{\sigma}^{-2}2^{t-1}\geq\underline{C}t^{3}\quad\forall\ t\geq 1,$
and note that $\eta_{t}^{-2}\geq\underline{\sigma}^{-2}2^{t-1}$ for all $\sigma<\underline{\sigma}$. The conclusion then follows from Lemma 8. ∎
A.3  Proposition 1
Proof.
Without loss of generality, we assume $\lambda_{0}>\frac{1}{2}$.
Parts (i) and (ii): First, we prove Part (ii). Fix $\Sigma\geq\Sigma^{\prime}$. Then it follows that $A_{t}:=\sqrt{\eta_{t}^{2}+\eta_{t-1}^{2}}\geq\sqrt{(\eta_{t}^{\prime})^{2}+(\eta_{t-1}^{\prime})^{2}}=:A_{t}^{\prime}$ for all $t\geq 2$. Let $(\gamma_{t})_{t}$ and $(\gamma_{t}^{\prime})_{t}$ be the induced sequences, respectively. We use induction to show that $\gamma_{t}\leq\gamma_{t}^{\prime}$ for all $t\geq 1$. For the base case, note that $\gamma_{1}=\Phi^{-1}(\lambda_{0})=\gamma_{1}^{\prime}$. For the inductive step, assume $\gamma_{t}\leq\gamma_{t}^{\prime}$. By Lemma 5,
$\gamma_{t+1}=F_{A_{t+1}}(\gamma_{t})\leq F_{A_{t+1}^{\prime}}(\gamma_{t})\leq F_{A_{t+1}^{\prime}}(\gamma_{t}^{\prime})=\gamma_{t+1}^{\prime}$
Hence, $\gamma_{\infty}\leq\gamma_{\infty}^{\prime}$ which implies
$\mu_{\infty}^{*}=c-\Phi(\gamma_{\infty})\geq c-\Phi(\gamma_{\infty}^{\prime})=(\mu_{\infty}^{*})^{\prime}$
as desired.
To prove Part (i), assume $\lambda_{0}>(1/2)$ and $\Sigma=\Sigma^{\prime}$. Then,
$\lambda_{0}\geq\lambda_{0}^{\prime}\implies\gamma_{1}\geq\gamma_{1}^{\prime}\implies\gamma_{2}=F_{A_{2}}(\gamma_{1})\geq F_{A_{2}}(\gamma_{1}^{\prime})=\gamma_{2}^{\prime}$
$\implies\gamma_{t}\geq\gamma_{t}^{\prime}\quad\forall\ t\implies\gamma_{\infty}\geq\gamma_{\infty}^{\prime}\implies\mu_{\infty}^{*}\leq(\mu_{\infty}^{*})^{\prime},$
where we use the same inductive argument as above.
Part (iii). Let $\eta_{t}^{2}$ and $\tilde{\eta}_{t}^{2}$ denote the time-$t$ posterior variances under $\Sigma$ and $\Sigma^{s\leftrightarrow s^{\prime}}$, respectively. Then, it immediately follows that $\eta_{t}^{2}\geq\tilde{\eta}_{t}^{2}$ holds for all $t$. Hence, by an analogous argument as in Part (i), we have $\mu_{t}^{*}(\Sigma,\lambda_{0})\geq\mu_{t}^{*}(\Sigma^{s\leftrightarrow s^{\prime}},\lambda_{0})$ holds for all $t$, which implies $\mu_{\infty}^{*}(\Sigma,\lambda_{0})\geq\mu_{\infty}^{*}(\Sigma^{s\leftrightarrow s^{\prime}},\lambda_{0})$.
A.4  Proposition 2
We prove Proposition 2 in two steps. First, we show that for any monotone selection rule with threshold $\mu^{*}\in(c-\lambda_{0},c-(1/2))$, there exists an exponential sequence $(A_{t})_{t}$ with $A_{t}\downarrow 0$ as $t\to\infty$ that implements it, i.e., $\mu_{\infty}^{*}=\mu^{*}$. Second, we show that there exists a noise process $\Sigma=(\sigma_{t})_{t}$ that implements the sequence $(A_{t})_{t}$ constructed in the first step.
Proof.
Without loss of generality, we assume $\lambda_{0}>\frac{1}{2}$.
Step 1. Fix any $\mu^{*}\in(c-\lambda_{0},c-(1/2))$. Our goal is to construct an exponential sequence $(A_{t})_{t}$ that satisfies $A_{t}\downarrow 0$ as $t\to\infty$, $A_{t+1}^{2}-A_{t+2}^{2}\leq A_{t}^{2}-A_{t+1}^{2}$ for $t\geq 2$, and $\mu_{\infty}^{*}=\mu^{*}$.
By taking $t\to\infty$ for ($\mu^{*}$-LoM), we see that
$\mu_{\infty}^{*}+\Phi(\gamma_{\infty})=c$
and hence
$\mu_{\infty}^{*}=\mu^{*}\iff\gamma_{\infty}=\gamma^{*}:=\Phi^{-1}(c-\mu^{*})\in(0,\gamma_{1})$
where $\gamma_{1}:=\Phi^{-1}(\lambda_{0})$. Hence, it suffices to find an exponential sequence $(A_{t})_{t}$ whose induced $(\gamma_{t})_{t}$ via ($\gamma$-LoM) satisfies $\gamma_{t}\downarrow\gamma^{*}$. By Lemma 7, we can choose $\gamma_{2}\in(\gamma^{*},\gamma_{1})$ close enough to $\gamma^{*}$ such that
$A_{2}:=A_{\gamma_{1}}(\gamma_{2})>\sqrt{2}A_{\gamma^{*}}(\gamma_{2})>0$
because i) $A_{\gamma_{1}}(\gamma^{*})=\bar{A}^{*}>0=A_{\gamma^{*}}(\gamma^{*})$; ii) $A_{\gamma^{*}}(\gamma_{2})>0$ for $\gamma_{2}\in(\gamma^{*},m)$; and iii) $A_{\gamma_{1}}(\cdot)$ and $A_{\gamma^{*}}(\cdot)$ are continuous. For all $t\geq 3$, we define
$A_{t}:=A_{\gamma^{*}}(\gamma_{t-1})=\frac{\Phi(\gamma_{t-1})-\Phi(\gamma_{t})}{\gamma_{t}}\quad\text{and}\quad\gamma_{t}:=\frac{\gamma^{*}+\gamma_{t-1}}{2},$
which implies
$\gamma_{t}=\frac{\gamma_{2}-\gamma^{*}}{2^{t-2}}+\gamma^{*}.$
By construction, $\gamma_{t}\downarrow\gamma^{*}$ and $A_{t}\downarrow 0$. It remains to show that $(A_{t})_{t}$ is exponential. Note that for all $t\geq 3$,
$\gamma_{t-1}-\gamma_{t}=\gamma_{t}-\gamma^{*}=4(\gamma_{2}-\gamma^{*})2^{-t}.$
Since $\varphi$ is decreasing on $(0,\infty)$, we have
$\varphi(\gamma_{2})(\gamma_{t-1}-\gamma_{t})\leq\varphi(\gamma_{t-1})(\gamma_{t-1}-\gamma_{t})\leq\Phi(\gamma_{t-1})-\Phi(\gamma_{t})$
$\leq\varphi(\gamma_{t})(\gamma_{t-1}-\gamma_{t})\leq\varphi(\gamma^{*})(\gamma_{t-1}-\gamma_{t}),$
where the second and the third inequalities follow from the mean value theorem. Hence,
$\underbrace{4\frac{\varphi(\gamma_{2})}{\gamma_{2}}(\gamma_{2}-\gamma^{*})}_{=:c}2^{-t}\leq\frac{\Phi(\gamma_{t-1})-\Phi(\gamma_{t})}{\gamma_{2}}\leq A_{t}$
$\leq\frac{\Phi(\gamma_{t-1})-\Phi(\gamma_{t})}{\gamma^{*}}\leq\underbrace{4\frac{\varphi(\gamma^{*})}{\gamma^{*}}(\gamma_{2}-\gamma^{*})}_{=:C}2^{-t}$
as desired, where the second and the third inequalities follow from $\gamma^{*}<\gamma_{t}\leq\gamma_{2}$ for all $t\geq 2$.
Finally, we want to show that the sequence $(A_{t})_{t}$ constructed above satisfies $A_{t+1}^{2}-A_{t+2}^{2}\leq A_{t}^{2}-A_{t+1}^{2}$ for $t\geq 2$ if we take $\gamma_{2}$ sufficiently close to $\gamma_{2}$. We defined the sequence of $A_{t}$ so that $A_{2}\geq\sqrt{2}A_{3}$ holds. Hence, we have
$A_{2}\geq\sqrt{2}A_{3}\implies A_{2}^{2}\geq 2A_{3}^{2}\geq 2A_{3}^{2}-A_{4}^{2}\implies A_{3}^{2}-A_{4}^{2}\leq A_{2}^{2}-A_{3}^{2}$
as desired. For $t\geq 3$, note that
$c2^{-t}\leq A_{t}\leq C2^{-t}$
$\implies\left(c^{2}-\frac{C^{2}}{4}\right)2^{-2t}\leq A_{t}^{2}-A_{t+1}^{2}\leq\left(C^{2}-\frac{c^{2}}{4}\right)2^{-2t}$
Thus, a sufficient condition for $A_{t+1}^{2}-A_{t+2}^{2}\leq A_{t}^{2}-A_{t+1}^{2}$ is
$\left(C^{2}-\frac{c^{2}}{4}\right)2^{-2(t+1)}\leq\left(c^{2}-\frac{C^{2}}{4}\right)2^{-2t}\Leftrightarrow\frac{C}{c}=\frac{\varphi(\gamma^{*})/\gamma^{*}}{\varphi(\gamma_{2})/\gamma_{2}}\leq\sqrt{\frac{17}{8}},$
which is guaranteed by choosing $\gamma_{2}$ close enough to $\gamma^{*}$.
Step 2. To complete the proof of Proposition 2, we show that there exists a noise process $\Sigma=(\sigma_{t})_{t}$ that implements the sequence $(A_{t})_{t}$ constructed above. First, we claim that it suffices to find a non-negative, weakly decreasing sequence $(\eta_{t}^{2})_{t}$ of posterior variances. This is true because given any such $(\eta_{t}^{2})_{t}$, the sequence $(\eta_{t}^{-2})_{t}$ is non-negative and weakly increasing. Thus, $(\eta_{t}^{2})_{t}$ is implemented by the noise process
$\sigma_{1}^{-2}:=\eta_{1}^{-2},\quad\sigma_{t}^{-2}=\eta_{t}^{-2}-\eta_{t-1}^{-2}\quad\forall\ t\geq 2$
Thus, it suffices to find a non-negative and weakly decreasing sequence $(\eta_{t}^{2})_{t\geq 1}$ such that
$A_{t}^{2}=\eta_{t}^{2}+\eta_{t-1}^{2}\quad\forall t\geq 2$
Note that any choice of $\eta_{1}^{2}$ pins down a candidate solution $(\eta_{t}^{2})_{t\geq 1}$ by $\eta_{t+1}^{2}=A_{t+1}^{2}-\eta_{t}^{2}$. We therefore proceed by stating the implied constraints on $\eta_{1}^{2}$.
Step 2.1. Let us begin by establishing the formula
$\eta_{t}^{2}(\eta_{1}^{2}):=(-1)^{t+1}\eta_{1}^{2}+\sum_{s=2}^{t}(-1)^{t+s}A_{s}^{2}$
This follows by induction. For the base case, note that $\eta_{2}^{2}=A_{2}^{2}-\eta_{1}^{2}$. For the inductive step, suppose the formula holds for $\eta_{t}^{2}$. Then,
$\eta_{t+1}^{2}(\eta_{1}^{2})=A_{t+1}^{2}-\eta_{t}^{2}(\eta_{1}^{2})$ $=A_{t+1}^{2}+(-1)^{(t+1)+1}\eta_{1}^{2}+\sum_{s=2}^{t}(-1)^{(t+1)+s}A_{s}^{2}$
$=(-1)^{t+2}\eta_{1}^{2}+\sum_{s=2}^{t+1}(-1)^{(t+1)+s}A_{s}^{2}$
since $2(t+1)$ is even.
Step 2.2. Next, we pin down the implied constraints on $\eta_{1}^{2}$. The base case constraints are:
$0\leq A_{2}^{2}-\eta_{1}^{2}\leq\eta_{1}^{2}\iff L_{1}:=\frac{A_{2}^{2}}{2}\leq\eta_{1}^{2}\leq A_{2}^{2}=:U_{1}$
For every $t\geq 2$, we have the constraint
$0\leq\eta_{t+1}^{2}\leq\eta_{t}^{2}$
We can split this set of constraints into two cases.
Case 1. If $t$ is even, then the time-$t$ constraint from above is
$\sum_{s=2}^{t+1}(-1)^{t+s}A_{s}^{2}\leq\eta_{1}^{2}\leq\sum_{s=2}^{t}(-1)^{t+s}A_{s}^{2}-\frac{A_{t+1}^{2}}{2}$
which is equivalent to
$L_{t}:=(A_{2}^{2}-A_{3}^{2})+\cdots+(A_{t}^{2}-A_{t+1}^{2})\leq\eta_{1}^{2}\leq(A_{2}^{2}-A_{3}^{2})+\cdots+(A_{t}^{2}-(A_{t+1}^{2}/2))=:U_{t}$
Case 2. If $t$ is odd, then the time-$t$ constraint from above is:
$\sum_{s=2}^{t}(-1)^{(t+1)+s}A_{s}^{2}+\frac{A_{t+1}^{2}}{2}\leq\eta_{1}^{2}\leq\sum_{s=2}^{t+1}(-1)^{(t+1)+s}A_{s}^{2}$
which is equivalent to:
$L_{t}:=(A_{2}^{2}-A_{3}^{2})+\cdots+(A_{t-1}^{2}-A_{t}^{2})+(A_{t+1}^{2}/2)\leq\eta_{1}^{2}$
$\leq(A_{2}^{2}-A_{3}^{2})+\cdots+(A_{t-1}^{2}-A_{t}^{2})+A_{t+1}^{2}=:U_{t}$
Note that for odd $t$, $L_{t}\geq L_{t-1}$ and for even $t$, $U_{t}\leq U_{t-1}$. Hence, only the lower bounds for odd $t$ and the upper bounds for even $t$ have bite.
Step 2.3. Finally, we show that the previously derived properties of $(A_{t})_{t}$ imply that there exists $\eta_{1}^{2}$ satisfying the above constraints, and hence pins down a solution $\Sigma$. First, note that $(L_{t})_{t\text{ odd}}$ are increasing, since for any odd $t\geq 1$,
$L_{t}\leq L_{t+2}\iff\frac{A_{t+1}^{2}}{2}\leq A_{t+1}^{2}-A_{t+2}^{2}+\frac{A_{t+3}^{2}}{2}$
$\iff A_{t+2}^{2}-A_{t+3}^{2}\leq A_{t+1}^{2}-A_{t+2}^{2}$
which holds by Step 1. An analogous argument implies that $(U_{t})_{t\text{ even}}$ are decreasing. Second, note that $A_{t}^{2}$ decreasing implies $L_{t}\leq U_{t+1}$ for any odd $t$. Hence, by the nested intervals theorem, we have:
$\sup_{t\text{ odd}}L_{t}=\eta_{1}^{2}=\sum_{s=1}^{\infty}(A_{2s}^{2}-A_{2s+1}^{2})=\inf_{t\text{ even}}U_{t}$
and this sum converges because
$\left(c^{2}-\frac{C^{2}}{4}\right)2^{-4s}\leq A_{2s}^{2}-A_{2s+1}^{2}\leq\left(C^{2}-\frac{c^{2}}{4}\right)2^{-4s}$
$\Rightarrow\frac{1}{15}\left(c^{2}-\frac{C^{2}}{4}\right)\leq\sum_{s=1}^{\infty}(A_{2s}^{2}-A_{2s+1}^{2})\leq\frac{1}{15}\left(C^{2}-\frac{c^{2}}{4}\right).$
Note that $\eta_{1}^{2}$ is positive because $C/c\leq\sqrt{17/8}$ implies $c^{2}-C^{2}/4>0$. ∎
A.5  Proposition 3
Proof.
Without loss of generality, we assume $\lambda_{0}>\frac{1}{2}$. Part (i). Fix any $\theta\in(c-\lambda_{0},c-(1/2))$. For any $\sigma>0$, define $T(\sigma)\in\mathbb{N}$ implicitly as $\mu_{T(\sigma)}^{*}\leq\theta<\mu_{T(\sigma)+1}^{*}$. Note that $T(\sigma)$ is well-defined because for any fixed $\sigma>0$, $\mu_{t}^{*}\uparrow c-(1/2)$.
First, we show that $T(\sigma)\to\infty$ as $\sigma\downarrow 0$. Notice that $T(\sigma)$ is weakly decreasing in $\sigma$ because for any $\sigma^{\prime}>\sigma$, $\theta\geq\mu^{*}_{T(\sigma^{\prime})}(\sigma^{\prime})>\mu^{*}_{T(\sigma^{\prime})}(\sigma)$ implies $T(\sigma)\geq T(\sigma^{\prime})$ by the definition of $T(\sigma)$. Hence, $T(\sigma)$ either converges to some constant or diverges to $\infty$ as $\sigma\downarrow 0$. Suppose for a contradiction that $\lim_{\sigma\downarrow 0}T(\sigma)=T$ holds for some $T>0$. Since $T(\sigma)\uparrow T$ as $\sigma\downarrow 0$, for any $\sigma$, we have
$\mu^{*}_{T+1}(\sigma)>\mu^{*}_{T(\sigma)+1}(\sigma)>\theta,$
which implies $\lim_{\sigma\downarrow 0}\mu^{*}_{T+1}(\sigma)\geq\theta$. This contradicts $\lim_{\sigma\downarrow 0}\mu^{*}_{T+1}(\sigma)=c-\lambda_{0}$.
Next, fix any $\alpha\in(0,1)$. For any $0\leq t\leq\alpha T(\sigma)$, we have:
$\theta-\mu_{t}^{*}>\mu_{T(\sigma)}^{*}-\mu_{t}^{*}$ $=\sum_{s=t}^{T(\sigma)-1}(\mu_{s+1}^{*}-\mu_{s}^{*})=\sum_{s=t}^{T(\sigma)-1}\sigma\sqrt{\frac{1}{s+1}+\frac{1}{s}}\Phi^{-1}(c-\mu_{s+1}^{*})$
$\implies\lambda_{t}$ $=\Phi\left(\frac{\sqrt{t}(\theta-\mu_{t}^{*})}{\sigma}\right)>\Phi\left(\sum_{s=t}^{T(\sigma)-1}\sqrt{\frac{t}{s+1}+\frac{t}{s}}\Phi^{-1}(c-\mu_{s+1}^{*})\right)$
$\geq\Phi\left((T(\sigma)-t)\sqrt{\frac{t}{T(\sigma)}+\frac{t}{T(\sigma)-1}}\Phi^{-1}(c-\theta)\right)$
$\geq\Phi\left((1-\alpha)T(\sigma)\sqrt{\frac{1}{T(\sigma)}+\frac{1}{T(\sigma)-1}}\Phi^{-1}(c-\theta)\right)\to 1\quad\text{as}\quad\sigma\downarrow 0.$
Since the lower bound on $\theta-\mu_{t}^{*}$ holds uniformly for $t\in[0,\alpha T(\sigma)]$, for any $\epsilon>0$, there exists $\underline{\sigma}_{\alpha}>0$ such that for all $\sigma<\underline{\sigma}_{\alpha}$ and all $t\in[0,\alpha T(\sigma)]$, $\lambda_{t}\geq 1-\epsilon$.
Finally, fix any $\beta\in(1,\overline{\beta})$, where
$\overline{\beta}:=c+(1/2)-\theta-\delta\quad\text{and}\quad\delta=(c-(1/2)-\theta)/2$
and note that, for any $t\in[\beta T(\sigma),\infty)$,
$\theta-\mu_{t}^{*}\leq\theta-\mu_{\beta T(\sigma)}^{*}$ $<\mu_{T(\sigma)+1}^{*}-\mu_{\beta T(\sigma)}^{*}=-\sum_{s=T(\sigma)+1}^{\beta T(\sigma)-1}\sigma\sqrt{\frac{1}{s+1}+\frac{1}{s}}\Phi^{-1}(c-\mu_{s+1}^{*})$
$\implies\lambda_{t}$ $<\Phi\left(-\sum_{s=T(\sigma)+1}^{\beta T(\sigma)-1}\sqrt{\frac{t}{s+1}+\frac{t}{s}}\Phi^{-1}(c-\mu_{s+1}^{*})\right)$
$<\Phi\left(-\Big{(}(\beta-1)T(\sigma)-1\Big{)}\sqrt{1+\frac{\beta T(\sigma)}{\beta T(\sigma)-1}}\Phi^{-1}(c-\mu_{\beta T(\sigma)}^{*})\right)$
$<\Phi\left(-\Big{(}(\beta-1)T(\sigma)-1\Big{)}\sqrt{1+\frac{\beta T(\sigma)}{\beta T(\sigma)-1}}\Phi^{-1}(c-\theta-(\overline{\beta}-1))\right)$
$\to 0\quad\text{as }\sigma\to 0$
where the inequality follows from
$\mu_{\beta T(\sigma)}^{*}-\mu_{T(\sigma)}$ $=\sum_{s=T(\sigma)}^{\beta T(\sigma)-1}(\mu_{s+1}^{*}-\mu_{s}^{*})\leq\sum_{s=T(\sigma)}^{\beta T(\sigma)-1}\frac{1}{s}$
$\leq(\beta-1)T(\sigma)\frac{1}{T(\sigma)}=\beta-1$
$\implies c-\mu_{\beta T(\sigma)}^{*}$ $\geq c-\theta-(\beta-1)>c-\theta-(\overline{\beta}-1)>\frac{1}{2}$
$\implies\Phi^{-1}(c-\mu_{\beta T(\sigma)}^{*})$ $>\Phi^{-1}(c-\theta-(\overline{\beta}-1))>0.$
Since this upper bound on $\theta-\mu_{t}^{*}$ holds uniformly for $t\in[\beta T(\sigma),\infty)$, there exists $\underline{\sigma}_{\beta}>0$ such that for all $\sigma<\underline{\sigma}_{\beta}$ and all $t\in[\beta T(\sigma),\infty)$, $\lambda_{t}<\epsilon$.
Finally, define $\underline{\sigma}:=\min\{\underline{\sigma}_{\alpha},\underline{\sigma}_{\beta}\}$.
Part (ii). For any $t\geq 1$, we have
$|\lambda_{t+1}-\lambda_{t}|$ $=\left|\Phi\bigg{(}\cfrac{\theta-\mu^{*}_{t+1}}{\sigma/\sqrt{t+1}}\bigg{)}-\Phi\bigg{(}\cfrac{\theta-\mu^{*}_{t}}{\sigma/\sqrt{t}}\bigg{)}\right|$
$\leq\frac{1}{\sqrt{2\pi}\sigma}\left|(\sqrt{t+1}-\sqrt{t})\theta-\sqrt{t+1}\mu_{t+1}^{*}+\sqrt{t}\mu_{t}^{*}\right|$
$\leq\frac{\sqrt{t+1}-\sqrt{t}}{\sqrt{2\pi}\sigma}|\theta|+\frac{\sqrt{t+1}\mu_{t+1}^{*}-\sqrt{t}\mu_{t}^{*}}{\sqrt{2\pi}\sigma}$
$=\frac{\sqrt{t+1}-\sqrt{t}}{\sqrt{2\pi}\sigma}|\theta|+\frac{\sqrt{t+1}}{\sqrt{2\pi}\sigma}(\mu_{t+1}^{*}-\mu_{t}^{*})+\frac{(\sqrt{t+1}-\sqrt{t})\mu_{t}^{*}}{\sqrt{2\pi}\sigma}$
$<\frac{\sqrt{t+1}-\sqrt{t}}{\sqrt{2\pi}\sigma}|\theta|+\frac{\sqrt{t+1}}{\sqrt{2\pi}\sigma t}+\frac{c(\sqrt{t+1}-\sqrt{t})}{\sqrt{2\pi}\sigma}$
$\leq\frac{1}{\sqrt{2\pi}\sigma}\left\{(\sqrt{2}-1)(|\theta|+c)+\sqrt{2}\right\},$
where the second line follows from $\Phi^{\prime}(\cdot)\leq 1/\sqrt{2\pi}$, and the fifth line follows from $\mu_{t}^{*}<c$ and $\mu_{t+1}^{*}-\mu_{t}^{*}<t^{-1}$. Hence, for any $\epsilon$, there exists $\overline{\sigma}>0$ such that $|\lambda_{t+1}-\lambda_{t}|<\epsilon$ holds for all $t\geq 1$ and $\sigma>\overline{\sigma}$. Note that this argument does not depend on the value of $\theta$ as long as $|\theta|<\infty$ holds.
Appendix B Signals about past play
In the main text, we claimed that our model is equivalent to one in which players receive noisy signals of past play. We now formalize this claim by considering a single modification to our model: at each time $t$, each player $i\in[0,1]$ observes the independent signals
$x_{it}\sim N\Big{(}\theta,\sigma^{2}_{t}\Big{)}\quad\text{and}\quad y_{it}\sim N\Big{(}\Phi^{-1}(\lambda_{t-1}),\tau^{2}_{t}\Big{)}\text{ for $t>1$}$
This formulation follows that of Dasgupta (2007) and Trevino (2020).
Proposition 4.
$(\lambda_{t})_{t}$ is consistent with some $\Sigma$ under our main model if and only if $(\lambda_{t})_{t}$ is consistent with some $(\tau_{t}^{2},\sigma_{t}^{2})_{t\geq 1}$ under the Appendix B model.
Proof.
Notice that the environment in the main text can be viewed as the case in which $\tau^{2}_{t}=+\infty$ for all $t\geq 0$. Also observe that the case in which $\sigma^{2}_{1}<+\infty$, $\sigma^{2}_{t}=+\infty$ for all $t>1$ can be viewed as a social learning environment: at the start of the game, each player receives an independent signal of the state and, over the course of the game, the path of play also aggregates information about the state.
We now show that the dynamics of the game induced by the path $(\tau^{2}_{t},\sigma^{2}_{t})_{t\geq 1}$ can be replicated by some alternate path $(\sigma^{\prime}{{}^{2}}_{t})_{t\geq 1}$ in our main model in which past play is unobserved. Let $\eta_{t}^{-2}$ and $\eta_{t}^{\prime-2}$ be the posterior precisions induced by $(\tau^{2}_{t},\sigma^{2}_{t})_{t\geq 1}$ and $\Sigma^{\prime}$, respectively.
As before, let $\eta^{2}_{t}$ be the posterior variance at time $t$ for $t\geq 2$ with the initial conditions $\mu_{i1}=x_{i1}$ and $\eta_{1}^{2}=\sigma_{1}^{2}$. Then, the law of motion is unchanged from the main model and given by Lemma 1.
We can show the law of motion of $\mu_{it}$ and $\eta_{t}^{2}$ and ($\lambda$-LoM) by induction. Define $I_{is}=(x_{is},y_{is})_{s\leq t}$. At $t=1$, agent $i$’s posterior is given by $\theta|I_{i1}\sim N(x_{i1},\sigma_{1}^{2})$, which implies $\mu_{i1}=x_{i1}$ and $\eta_{1}^{2}=\sigma_{1}^{2}$. Since $\lambda_{0}$ is common knowledge, she takes $a_{i1}=1$ if and only if $\mu_{i1}\geq c-\lambda_{0}=\mu_{1}^{*}$. Since $\mu_{i1}|\theta\sim N(\theta,\eta_{1}^{2})$, $\lambda_{1}$ is given by
$\lambda_{1}=\text{Pr}(x_{i1}\geq\mu_{1}^{*})=\Phi\left(\frac{\theta-\mu_{1}^{*}}{\sigma}\right).$
Now suppose that the statement holds for $t-1$. By assumption, we have $\Phi^{-1}(\lambda_{t-1})=(\theta-\mu_{t-1}^{*})/\eta_{t-1}$, which implies
$\mu_{it-1}\,|\,\theta$ $\sim N(\theta,\eta_{t-1}^{2}),$
$x_{it}\,|\,\theta$ $\sim N(\theta,\sigma_{t}^{2}),$
$(\eta_{t-1}y_{it}+\mu_{t-1}^{*})\,|\,\theta$ $\sim N(\theta,\tau_{t}^{2}\eta_{t-1}^{2}).$
Since they are independent of each other conditional on $\theta$, Bayesian updating yields $\eta_{1}^{-2}=\sigma_{1}^{-2}$, $\mu_{i1}=x_{i1}$, and for each $t\geq 2$,
$\eta_{t}^{-2}=\eta_{t-1}^{-2}(1+\tau_{t}^{-2})+\sigma_{t}^{-2}$
$\mu_{it}=\sum_{s=1}^{t}\left(\frac{\eta_{s-1}^{-2}\tau_{s}^{-2}}{\eta_{t}^{-2}}\tilde{y}_{is}+\frac{\sigma_{s}^{-2}}{\eta_{t}^{-2}}x_{is}\right)$
where $\tau_{1}^{-2}:=0$. This implies:
$\eta_{t}^{-2}=\sum_{s=1}^{t}\sigma_{s}^{-2}\prod_{r>s}^{t}(1+\tau_{r}^{-2})$
Since ($\mu^{*}$-LoM) remains unchanged from Lemma 1, we obtain ($\lambda$-LoM) by the proof of Lemma 1. Then, we can find a sequence $(\sigma^{\prime}_{t})_{t}$ such that the dynamics are equivalent:
$\sigma_{t}^{\prime-2}:=\eta_{t}^{-2}-\eta_{t-1}^{-2}$
Appendix C Finite Player Approximation
Consider the following game, which we denote as the $N$-player model, in contrast to the main text’s continuum model. There are $N$ players, indexed by $i\in[N]$. Actions, timing, and information are the same as in the main text. [^30] Define the random variables $(x_{it})_{i\geq 1,t\geq 1}$ on the same probability space. Agent $i$’s payoffs at time $t$ are:
$u(a_{it},\lambda_{N,-i,t-1},\theta):=\begin{cases}\theta+\lambda_{N,-i,t-1}&\text{if }a_{it}=1\\ c&\text{if }a_{it}=0\\ \end{cases}$
where
$\lambda_{N,-i,t-1}=\frac{1}{N-1}\sum_{j\in[N]\setminus\{i\}}\bm{1}[a_{j,t-1}=1]$
is the empirical proportion of non-$i$ agents who took the risky action at time $t-1$ for $t>1$, and $\lambda_{N,-i,t-1}:=\lambda_{0}$ is common knowledge. Let
$\lambda_{N,t}:=\frac{1}{N}\sum_{i=1}^{N}\bm{1}[a_{it}=1]$
be the empirical proportion of all agents who attack at time $t$.
The purpose of Appendix C is to show that the dynamics of the $N$-player model converge to the dynamics of the continuum model as $N$ becomes large. Let us briefly discuss this exercise. With a finite number of players, there is randomness in the empirical distribution of beliefs at time-$t$. This, in turn, translates into randomness in the path of aggregate play. Theorem 2 shows that in an environment with many players, this randomness washes out.
To show Theorem 2, we begin by analyzing the dynamics of belief thresholds in the $N$-player model.
Lemma 9.
At each time $t\geq 1$, each agent $i\in[N]$ takes the risky action if and only if $\mu_{it}\geq\mu_{t}^{*}$, where $(\mu_{t}^{*})_{t\geq 1}$ is the same sequence of thresholds as in the main model.
Proof.
Base case: At time $t=1$, each agent $i$ takes the risky action iff $\mu_{i1}=x_{i1}\geq c-\lambda_{0}=:\mu_{1}^{*}$.
Inductive step: Fix $t>1$ and assume that at time $t-1$, each agent $i$ takes the risky action iff $\mu_{i,t-1}\geq\mu_{t-1}^{*}$. At time $t$, each agent $i$ takes the risky action iff
$\mathbb{E}_{it}\Big{[}\theta+\lambda_{N,i,t-1}\Big{]}=\mu_{it}+\frac{1}{N-1}\sum_{j\neq i}\mathbb{P}_{it}\Big{(}\mu_{j,t-1}\geq\mu_{t-1}^{*}\Big{)}\geq c$
Note that at time $t$, agent $i$ believes $\theta\sim N(\mu_{it},\eta_{t}^{2})$ and for each $j\neq i$, $\mu_{j,t-1}|\theta\sim N(\theta,\eta_{t-1}^{2})$. Hence, at time $t$ agent $i$ believes
$\mu_{j,t-1}\sim N(\mu_{it},\eta_{t-1}^{2}+\eta_{t}^{2})$
which implies she takes the risky action at time $t$ iff
$\mu_{it}+\Phi\left(\frac{\mu_{it}-\mu_{t-1}^{*}}{\sqrt{\eta_{t-1}^{2}+\eta_{t}^{2}}}\right)\geq c\iff\mu_{it}\geq\mu_{t}^{*}$
as desired. ∎
With this lemma in hand, observe that
$\lambda_{N,t}=1-F_{N,t}(\mu_{t}^{*})\quad\text{and}\quad\lambda_{t}(\theta)=1-F_{\theta,t}(\mu_{t}^{*})$
where $F_{N,t}$ is the empirical CDF of $N$ draws of $\mu_{it}$ and $F_{\theta,t}$ is the CDF of $\mu_{it}\sim N(\theta,\eta_{t}^{2})$.
Theorem 2.
For any $(\Sigma,\lambda_{0})$ and $\theta\in\mathbb{R}$,
[…]
Proof.
For each $N\geq 1$ and each $i\in[N]$, note that $\mu_{i,\infty}=\theta$ $\mathbb{P}_{\theta}$-a.s. by Doob’s theorem. Hence, $\lambda_{N,\infty}=\lambda_{\infty}(\theta)$ $\mathbb{P}_{\theta}$-a.s. ∎
Appendix D Alternative Interpretations of Model
Overlapping generations model. There is an initial unit mass of agents with proportion $\lambda_{0}\in(0,1)$ who take the risky action, all of whom share a common improper uniform prior about $\theta$. At each time $t\geq 1$, a unit mass of children is born (generation $t$), each of whom inherits her prior about $\theta$ from her parent’s (generation $t-1$) belief. Each child $i$ receives a signal $x_{it}\sim N(\theta,\sigma_{t}^{2})$, updates her belief about $\theta$, chooses an action, and gets uniformly matched to play the static game below with a member of generation $t-1$. Finally, child $i$ becomes parent $i$ and continues to take her time-$t$ action whenever matched with an agent from generation $t+1$. We interpret this as due to switching costs—since agents’ actions when playing against the future generation are not payoff-relevant, switching costs ensure this.
Two-player interpretation. Consider the set of static symmetric two-player games, indexed by the state $\theta\in\mathbb{R}$, whose payoff matrix is
[…]
Note that for $\theta\in(c-1,c)$, there are two pure strategy equilibria: $(A,A)$ and $(N,N)$. Applying Harsanyi, Selten et al. (1988)’s risk dominance criterion selects $(A,A)$ if and only if $\theta\geq c-(1/2)$.
Carlsson and Van Damme (1993) propose a static global games framework in which two players play the incomplete information game with state-dependent matrix from above. [^31] Our static two-player game corresponds to Figure 1 of Carlsson and Van Damme (1993) with $x=4(c-\theta)$, $\alpha_{1}=\alpha_{2}=N$, and $\beta_{1}=\beta_{2}=A$. Consider the case where both players share a common improper uniform prior, and each receive iid Gaussian signals distributed $N(\theta,\sigma^{2})$. Carlsson and Van Damme (1993) then show that as $\sigma\downarrow 0$, the risk dominant equilibrium is selected.
In this section, we show a tight connection between the above static global games setting and our model where $\Sigma=(\sigma^{2},\infty,\infty,\ldots)$, which we study in Example 1. In this case, $(\mu_{t}^{*})_{t}$ evolves according to the law of motion
$\mu_{t}^{*}+\Phi\left(\frac{\mu_{t}^{*}-\mu_{t-1}^{*}}{\sqrt{2}\sigma}\right)=c$
with $\mu_{1}^{*}=c-\lambda_{0}$. As we have shown, this equation has a unique fixed point at $c-(1/2)$ and hence we have global convergence $\mu_{t}^{*}\to c-(1/2)$ for any $\lambda_{0}\in(0,1)$. Furthermore, the entire path of $(\mu_{t}^{*})_{t}$ corresponds to the iterated best-response dynamics with symmetric switching strategies in Carlsson and Van Damme (1993)’s two-player setting with the informational environment from above, in which $\mu_{\infty}^{*}$ is the unique rationalizable switching strategy. To see this, let $\mathbb{P}_{x}(\theta,x^{\prime})$ denote the joint posterior over the state $\theta$ and my opponent’s signal $x^{\prime}$, conditional on my signal $x$.
The key observation is that, if my opponent follows a switching strategy with cutoff $\mu^{\prime}$, my expected payoff from $A$ given $x$ is
$\mathbb{P}_{x}(x^{\prime}\geq\mu^{\prime})\Big{(}\mathbb{E}_{x}[\theta|x^{\prime}\geq\mu^{\prime}]+1\Big{)}+\mathbb{P}_{x}(x^{\prime}<\mu^{\prime})\mathbb{E}_{x}[\theta|x^{\prime}<\mu^{\prime}]$
[…]
which proves that if my opponent follows a switching strategy with cutoff $\mu_{t-1}^{*}$, my best response is to follow a switching strategy with cutoff $\mu_{t}^{*}$, as desired.
An analogous argument allows us to extend the interpretation of our model dynamics as two-player iterated joint learning and best-response dynamics by considering our model for arbitrary $\Sigma$. In this case, note that if my opponent receives signal distributed $N(\theta,\eta_{t-1}^{2})$ (corresponding to the time-$(t-1)$ distribution of posterior means, conditional on the state) and my belief is about the state is $N(\mu_{it},\eta_{t}^{2})$, then my (state-unconditional) belief about my opponent’s signal is
$x^{\prime}\sim N(\mu_{it},\eta_{t-1}^{2}+\eta_{t}^{2})\implies\mathbb{P}_{it}(x^{\prime}\geq\mu^{\prime})=\Phi\left(\frac{\mu_{it}-\mu^{\prime}}{\sqrt{\eta_{t-1}^{2}+\eta_{t}^{2}}}\right)$
and hence our general setting is also consistent with iterated joint learning and best-response dynamics. This can be interpreted as best-response dynamics with bounded rationality, since at each time each player is only able to take into account one extra signal.
Appendix E Payoff extensions
In the main text, we assumed that the payoff of taking action $1$ is the sum of the state and the measure of agents who played action $1$ in the previous period. Now we relax this assumption and suppose agent $i$’s payoffs at time $t$ are:
$u(a_{it},\lambda_{t-1},\theta):=\begin{cases}\alpha\theta+(1-\alpha)\lambda_{t-1}&\text{if }a_{it}=1\\ c&\text{if }a_{it}=0\\ \end{cases}$
with $\alpha\in(0,1).$ The parameter of $\alpha$ can be interpreted as the inverse of the extent of intertemporal strategic complementarity. To simplify the notations, we define the followings:
$\tilde{\theta}(\alpha)$ $:=\frac{\alpha}{1-\alpha}\theta$
$\tilde{c}(\alpha)$ $:=\frac{c}{1-\alpha}$
$\tilde{\sigma}_{t}^{2}(\alpha)$ $:=\left(\frac{\alpha}{1-\alpha}\right)^{2}\sigma_{t}^{2}$
$\tilde{\eta}_{t}^{2}(\alpha)$ $:=\left(\frac{\alpha}{1-\alpha}\right)^{2}\eta_{t}^{2}.$
In this environment, we define the risk dominant action at state $\theta$ as
$RD(\theta):=\mathbbm{1}\left[\theta\geq\frac{c}{\alpha}-\frac{1}{2}\frac{1-\alpha}{\alpha}\right]=\mathbbm{1}[\tilde{\theta}(\alpha)\geq\tilde{c}(\alpha)-(1/2)]$
i.e., the action that one would take if she knew the state and conjectured that half of all agents take the risky action. Analogously, we define the non-risk dominant action as $NRD(\theta):=1-RD(\theta)$.
Proposition 5.
All the results in the main text hold by replacing $c$, $\sigma_{t}^{2}$, $\eta_{t}^{2}$, and $\bar{\beta}=1+|c-(1/2)-\theta|/2$ with $\tilde{c}(\alpha)$, $\tilde{\sigma}_{t}^{2}(\alpha)$, $\tilde{\eta}_{t}^{2}(\alpha)$, and $\tilde{\beta}:=1+\big{|}\tilde{c}(\alpha)-(1/2)-\tilde{\theta}(\alpha)\big{|}/2$, respectively.
Proof.
We show ($\mu^{*}$-LoM) and ($\lambda$-LoM) by induction:
$\mu_{t}^{*}+\Phi\left(\frac{\mu_{t}^{*}-\mu_{t-1}^{*}}{\sqrt{\tilde{\eta}_{t-1}^{2}(\alpha)+\tilde{\eta}_{t}^{2}(\alpha)}}\right)=\tilde{c}(\alpha)$
$\lambda_{t}(\theta)=\mathbbm{1}\left\{\tilde{\theta}(\alpha)\geq\mu_{t}^{*}\right\}=\Phi\left(\frac{\tilde{\theta}(\alpha)-\mu_{t}^{*}}{\tilde{\eta}_{t}(\alpha)}\right)$
At $t=1$, agent $i$’s posterior belief is given by $\theta|I_{i1}\sim N(x_{i1},\sigma_{1}^{2})$, which implies $\tilde{\theta}(\alpha)|I_{i1}\sim N(\tilde{\mu}_{i1},\tilde{\sigma}_{1}^{2}(\alpha))$ with $\tilde{\mu}_{i1}:=\mathbb{E}_{i1}[\tilde{\theta}(\alpha)]=\alpha x_{i1}/(1-\alpha)=:\tilde{x}_{i1}$. Since $\lambda_{0}$ is common knowledge, she takes $a_{i1}=1$ if and only if $\tilde{\mu}_{i1}\geq\tilde{c}-\lambda_{0}=\mu_{1}^{*}$. Since $\tilde{\mu}_{i1}|\tilde{\theta}(\alpha)\sim N(\tilde{\theta}(\alpha),\tilde{\sigma}_{1}^{2}(\alpha))$, $\lambda_{1}$ is given by
$\lambda_{1}=\mathbb{P}_{\theta}(\tilde{\mu}_{i1}\geq\mu_{1}^{*})=1-\Phi\left(\frac{\mu_{1}^{*}-\tilde{\theta}(\alpha)}{\tilde{\sigma}_{1}(\alpha)}\right)=\Phi\left(\frac{\tilde{\theta}(\alpha)-\mu_{1}^{*}}{\tilde{\sigma}_{1}(\alpha)}\right).$
Hence, ($\lambda$-LoM) holds for $t=1$.
Now, for $t>1$, suppose that $a_{j,t-1}=1$ if and only if $\tilde{\mu}_{j,t-1}\geq\mu_{t-1}^{*}$. By standard Gaussian-Gaussian Bayesian updating, we have $\tilde{\theta}(\alpha)|I_{it}\sim N(\tilde{\mu}_{it},\tilde{\eta}_{t}^{2}(\alpha))$, where
$\tilde{\mu}_{it}$ $=\frac{\tilde{\eta}_{t-1}^{-2}(\alpha)}{\tilde{\eta}_{t-1}^{-2}(\alpha)+\tilde{\sigma}_{t}^{-2}(\alpha)}\tilde{\mu}_{i,t-1}+\frac{\tilde{\sigma}_{t}^{-2}(\alpha)}{\tilde{\eta}_{t-1}^{-2}(\alpha)+\tilde{\sigma}_{t}^{-2}(\alpha)}\tilde{x}_{it}=\tilde{\eta}_{t}^{2}(\alpha)\sum_{s=1}^{t}\tilde{\sigma}_{s}^{-2}(\alpha)\tilde{x}_{is}$
$\tilde{\eta}_{t}^{2}(\alpha)$ $=\frac{\tilde{\eta}_{t-1}^{2}(\alpha)}{1+\tilde{\sigma}_{t}^{-2}(\alpha)\tilde{\eta}_{t-1}^{2}(\alpha)}=\left(\sum_{s=1}^{t}\tilde{\sigma}_{s}^{-2}(\alpha)\right)^{-1},$
where $\tilde{x}_{is}:=\alpha x_{is}/(1-\alpha)$. Then, by an analogous argument as in the proof of Lemma 1, ($\mu^{*}$-LoM) and ($\lambda$-LoM) hold for $t$. Once we obtain ($\mu^{*}$-LoM) and ($\lambda$-LoM), the rest of the arguments carry through by replacing $c$, $\sigma_{t}^{2}$, $\eta_{t}^{2}$, and $\bar{\beta}=1+|c-(1/2)-\theta|/2$ with $\tilde{c}(\alpha)$, $\tilde{\sigma}_{t}^{2}(\alpha)$, $\tilde{\eta}_{t}^{2}(\alpha)$, and $\tilde{\beta}:=1+\big{|}\tilde{c}(\alpha)-(1/2)-\tilde{\theta}(\alpha)\big{|}/2$, respectively. ∎
In particular, since learning rates are unaffected by scaling, the tight connection between learning rates and limit play in Theorem 1 is robust to the extent of coordination:
Corollary 3.
Theorem 1 holds exactly as stated in the main text, without replacing the above expressions with their $\alpha$-counterparts.
Proof.
Note that $\eta_{t}^{-2}=O(t^{p})$ if and only if $\tilde{\eta}_{t}^{-2}(\alpha)=O(t^{p})$, and that $\eta_{t}^{-2}=\Omega(t^{q})$ if and only if $\tilde{\eta}_{t}^{-2}(\alpha)=\Omega(t^{q})$. Hence, Theorem 1 holds without replacing the above expressions with their $\alpha$-counterparts. ∎